### v2

# Kernel-Based Nonparametric Tests For Shape Constraints

Abstract.
We develop a reproducing kernel Hilbert space (RKHS) framework for nonparametric mean-variance optimization and inference on shape constraints of the optimal rule. We derive statistical properties of the sample estimator and provide rigorous theoretical guarantees, such as asymptotic consistency, a functional central limit theorem, and a finite-sample deviation bound that matches the Monte Carlo rate up to regularization. Building on these findings, we introduce a joint Wald-type statistic to test for shape constraints over finite grids. The approach comes with an efficient computational procedure based on a pivoted Cholesky factorization, facilitating scalability to large datasets. Empirical tests suggest favorably of the proposed methodology.
The author is grateful to Paul Schneider, Michael Multerer, and Emanuele Luzzi for helpful comments.
MSC 2020: 62G10, 62G20, 62P05, 46E22; JEL: C13, C14.
Keywords: Kernel methods, nonparametric estimation, shape-constrained inference, positivity/monotonicty/convexity tests, mean-variance optimization.
1. Introduction
Many modern learning and decision problems require not only accurate prediction of the level of an unknown function but also reliable control of its local behavior-positivity, monotonicity, convexity, and other shape features that are naturally expressed through derivatives. Applications where shape constraints play an important role include economics and asset pricing (Rochet and Choné (1998); Linn et al. (2017); Rosenberg and Engle (2002); Ait-Sahalia and Lo (2000); Jackwerth (2015)), optimal transport problems (Makkuva et al. (2020)), to name a few. In risk-sensitive tasks, it is often desirable to optimize a concave performance functional that depends on the value and the derivatives of an unknown function, while simultaneously quantifying uncertainty in those derivative functionals. While many works estimate functions under shape constraints, fewer provide formal statistical tests to assess whether such constraints hold in the population, particularly in flexible RKHS settings. This work develops a non-parametric framework based on reproducing kernel Hilbert space (RKHS) for testing constraints by embedding sufficiently smooth regression functions in the RKHS. This treats derivative evaluations as bounded linear functionals via derivative reproducing properties, see Zhou (2008), and also allows for a rigorous analysis of both asymptotic and finite-sample behavior of the optimal sample estimator. Embedding the learning problem in an RKHS not only ensures computational tractability through representer theorems but also allows treating derivative evaluations as bounded linear functionals, enabling a unified treatment of function and shape estimation.
1.1. Related work
Our work builds upon the following lines of research. The first is concerned with shape-constrained regression tasks, wherein the estimator is restricted to be a positive/monotone/convex function, see Groeneboom and Jongbloed (2014); Seijo and Sen (2011); Marteau-Ferey et al. (2020); Muzellec et al. (2022); Aubin-Frankowski and Szabo (2022) and references therein. The second line of research focuses on the estimation of these shape restrictions in regression problems, which is more aligned with our work, see Silvapulle and Sen (2001) and references therein. Parametric tests for econometric models have been developed in Shapiro (1985); Wolak (1987, 1989); Andrews (1998) to name a few, while nonparametric tests for shape constraints have been investigated in Ghosal et al. (2000); Hall and Heckman (2000); Juditsky and Nemirovski (2002); Birke and Neumeyer (2013). The non-parametric tests have been designed mostly for local averaging using kernel smoothing techniques, but do not discuss computational scalability. Yet another relevant field is learning theory, where asymptotic and finite-sample statistical properties have been derived in an RKHS framework, along with representer theorems. A few examples include Schölkopf et al. (2001); Cucker and Smale (2001); Cristianini and Schölkopf (2002); Caponnetto and De Vito (2007); Alaoui and Mahoney (2015); Filipović and Schneider (2025). However, most of these works are developed purely in the context of statistical learning and do not address the question of shape constraints in the estimation problem. Unlike shape-restricted estimation that enforces constraints during fitting, we estimate an unconstrained RKHS rule and test directional shape via finite-dimensional cone projections of derivative evaluations with plug-in covariance arising from a mean–variance objective.
1.2. Contributions
Our contributions are as follows:
  1. (1)
    We formulate a general mean-variance learning problem in RKHS built from a linear functional of function values and their gradients up to a fixed order. We establish the characterizations of the population and empirical optimizers, and derive a representer theorem that reduces the computation of the optimal empirical solution to a finite-dimensional system of equations.
  2. (2)
    We derive rigorous statistical guarantees for the empirical optimizer, including consistency, a functional central limit theorem (implying asymptotic normality of derivative evaluations), and finite-sample deviation bounds depending on sample size and regularization.
  3. (3)
    We propose a Wald-type test statistic for assessing shape restrictions over a finite grid (positivity, monotonicity, convexity, etc.). The test statistic measures a squared Mahalanobis distance of the projection error admits an implementation based on a non-negative least squares program.
  4. (4)
    We provide an efficient computation procedure based on a pivoted Cholesky decomposition that can be handle large datasets, with large samples, and can be used for testing on dense grids.
1.3. Outline
The remainder of this article is organized as follows. In Section 2, we set up our problem in an RKHS, and derive characterizations of the optimal solutions to the population and empirical problems; additionally, we state and prove the representer theorem. In Section 3, we derive the statistical properties of the sample estimator, including consistency and asymptotic distribution, as well as finite-sample error bounds. In Section 4, we construct the test statistic and detail the steps for inference on shape constraints. Section 5 addresses numerical experiments, where we showcase the efficacy of the developed methodology. In Section 6, we conclude and identify areas for future research.
2. Preliminaries
In this section, we first fix the notation and recall certain facts about reproducing kernel Hilbert spaces and operators on Hilbert spaces that will be useful for the remainder of the paper. We refer an interested reader to Reed and Simon (1981); Schatten (1970); Dunford and Schwartz (1958) for further details on the following.
2.1. Notation and setting
Let $\mathcal{H}$ be a separable Hilbert space with orthonormal basis $(e_{j})_{j\in\mathbb{N}}$. For a linear operator $A\colon\mathcal{H}\to\mathcal{H}$, denote its adjoint by $A^{\ast}$ and set $|A|\mathrel{\mathrel{\mathop{:}}=}(A^{\ast}A)^{1/2}$. We denote by $\mathscr{B}(\mathcal{H})$ the Banach space of bounded linear operators on $\mathcal{H}$ with the operator norm $\|A\|_{\text{op}}\mathrel{\mathrel{\mathop{:}}=}\sup\{\|Af\|_{\mathcal{H}}:f\in\mathcal{H},\|f\|_{\mathcal{H}}\leq 1\}$. The space of Hilbert-Schmidt operators $\mathscr{H}\mathscr{S}(\mathcal{H})\mathrel{\mathrel{\mathop{:}}=}\{A\in\mathscr{B}(\mathcal{H}):\|A\|_{\operatorname{HS}}<\infty\}$ is a Hilbert space, equipped with the inner product $\langle A,B\rangle_{\operatorname{HS}}\mathrel{\mathrel{\mathop{:}}=}\text{tr}(B^{\ast}A)=\sum_{j\in\mathbb{N}}\langle Ae_{j},Be_{j}\rangle_{\mathcal{H}}$, the sum being independent of the orthonormal basis. The space of trace-class (nuclear) operators defined as $\mathscr{T}(\mathcal{H})\mathrel{\mathrel{\mathop{:}}=}\{A\in\mathscr{B}(\mathcal{H}):\text{tr}(|A|)<\infty\}$ is a Banach space. Furthermore, we have the continuous inclusions $\mathscr{T}(\mathcal{H})\subset\mathscr{H}\mathscr{S}(\mathcal{H})\subset\mathscr{B}(\mathcal{H})$ with the norm bounds $\|A\|_{\text{op}}\leq\|A\|_{\operatorname{HS}}\leq\text{tr}(|A|)$. In particular, if $A\geq 0$, then $\text{tr}(|A|)=\text{tr}(A)$. For $f,g\in\mathcal{H}$, the rank-one operator $f\otimes g\colon\mathcal{H}\to\mathcal{H}$ is $(f\otimes g)u\mathrel{\mathrel{\mathop{:}}=}\langle u,g\rangle_{\mathcal{H}}\,f$. It satisfies $\langle A,\,f\otimes g\rangle_{\operatorname{HS}}=\langle Ag,\,f\rangle_{\mathcal{H}}$ and $\|f\otimes g\|_{\operatorname{HS}}=\|f\|_{\mathcal{H}}\,\|g\|_{\mathcal{H}}$. If $A$ is self-adjoint, we write $\sigma(A)$ for its spectrum and set $\lambda_{\min}(A)\mathrel{\mathrel{\mathop{:}}=}\inf\sigma(A)$, $\lambda_{\max}(A)\mathrel{\mathrel{\mathop{:}}=}\sup\sigma(A)$. For $s\in\mathbb{N}$, $\mathscr{C}^{s}(\mathcal{X})$ denotes the set of $s$ times continuously differentiable functions from $\mathcal{X}$ to $\mathbb{R}$. For a function $f$ of $d$ variables, and any multi-index $\boldsymbol{\alpha}\mathrel{\mathrel{\mathop{:}}=}(\alpha_{1},\ldots,\alpha_{d})\in\mathbb{N}^{d}$ with $|\boldsymbol{\alpha}|\mathrel{\mathrel{\mathop{:}}=}\alpha_{1}+\cdots+\alpha_{d}\leq s$, we denote the corresponding partial derivative of $f$ (when it exists),
$D^{\boldsymbol{\alpha}}f(\boldsymbol{x})\mathrel{\mathrel{\mathop{:}}=}\frac{\partial^{|\boldsymbol{\alpha}|}}{\partial x_{1}^{\alpha_{1}}\ldots\partial x_{d}^{\alpha_{d}}}f(\boldsymbol{x}).$
We define the set $\mathcal{A}_{s}\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{\alpha}\in\mathbb{N}^{d}:|\boldsymbol{\alpha}|\leq s\}$ and $m_{s}\mathrel{\mathrel{\mathop{:}}=}|\mathcal{A}_{s}|=\binom{s+d}{d}$. In addition, we utilize the usual notions of $o_{\mathbb{P}}$ and $\mathcal{O}_{\mathbb{P}}$, and refer the reader to van der Vaart (1998) for details.
2.2. Reproducing kernel Hilbert spaces
We recap a fundamental notion in statistical machine learning, namely that of a reproducing kernel Hilbert space. For more background and applications, we refer the reader to Wendland (2005); Berlinet and Thomas-Agnan (2004); Hastie et al. (2001). Let $\mathcal{X}\subset\mathbb{R}^{d}$ and let $\mathcal{K}\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ be a symmetric function such that for any finite set $\{\boldsymbol{x}_{1},\ldots,\boldsymbol{x}_{N}\}\subset\mathcal{X}$, the Gram matrix $\boldsymbol{K}\mathrel{\mathrel{\mathop{:}}=}\left[\mathcal{K}(\boldsymbol{x}_{i},\boldsymbol{x}_{j})\right]_{i,j=1}^{N}\in\mathbb{R}^{N\times N}$ is symmetric and positive semidefinite. The RKHS $\mathcal{H}$ associated with the kernel function $\mathcal{K}$ is defined to the completion of $\operatorname{span}\{\phi(\boldsymbol{x})\mathrel{\mathrel{\mathop{:}}=}\mathcal{K}(\boldsymbol{x},\cdot):\boldsymbol{x}\in\mathcal{X}\}$ with the inner product $\langle\cdot,\cdot\rangle_{\mathcal{H}}$ given by $\langle\phi(\boldsymbol{x}),\phi(\boldsymbol{y})\rangle_{\mathcal{H}}=\mathcal{K}(\boldsymbol{x},\boldsymbol{y})$. In this case, $\phi(\boldsymbol{x})$ acts as the unique Riesz representer of the evaluation functional at $\boldsymbol{x}\in\mathcal{X}$, and we call $\mathcal{K}$ the reproducing kernel of $\mathcal{H}$. The reproducing property says that
(2.1) $f(\boldsymbol{x})=\langle f,\phi(\boldsymbol{x})\rangle_{\mathcal{H}}\quad\text{for any }f\in\mathcal{H},\,\boldsymbol{x}\in\mathcal{X}.$
For a sufficiently smooth kernel $\mathcal{K}$ on any separable $\mathcal{X}$, the RKHS $\mathcal{H}$ is separable, see Cristianini and Schölkopf (2002, Lemma 4.3). Furthermore, we have the following result.
Theorem 2.1 (Zhou (2008, Theorem 1)).
Let $s\in\mathbb{N}$ and $\mathcal{K}\colon\mathcal{X}\times\mathcal{X}\to\mathbb{R}$ is a reproducing kernel such that $\mathcal{K}\in\mathscr{C}^{2s}(\mathcal{X}\times\mathcal{X})$. Then, it holds,
  • for any $\boldsymbol{x}\in\mathcal{X}$ and $\boldsymbol{\alpha}\in\mathcal{A}_{s}$, it holds, $\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x})\in\mathcal{H}$, where $\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x})\mathrel{\mathrel{\mathop{:}}=}D^{\boldsymbol{\alpha}}\mathcal{K}(\boldsymbol{x},\cdot)$;
  • a reproducing property holds for the partial derivatives for any $\boldsymbol{\alpha}\in\mathcal{A}_{s}$:
    (2.2) $D^{\boldsymbol{\alpha}}f(\boldsymbol{x})=\langle f,\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x})\rangle_{\mathcal{H}}\quad\text{for any }f\in\mathcal{H},\,\boldsymbol{x}\in\mathcal{X}.$
2.3. Mean-variance optimization in RKHS
One of the key advantages of formulating learning problems in an RKHS framework lies in the use of representer theorems; these facilitate a finite-dimensional formulation of the problem at hand that may be solved with conventional linear algebra techniques. Our result, see Theorem 2.4, is a variant thereof. Moreover, many learning problems involve the use of gradient information for better learning ability. For sufficiently smooth kernels, derivatives can be interpreted as bounded linear functionals in the RKHS via the reproducing property, which enables us to model nonparametrically the shape constraints in learning problems.
A related class of problems seeks to maximize a concave utility of a task-specific functional that depends on the value and derivative information of an unknown underlying function. In this setting, a mean-variance objective provides a principled way to balance the expected performance with variability, thereby capturing risk awareness and down-weighting high-uncertainty regions. For example, one may optimize a portfolio decision rule whose payoff depends on an underlying function and its gradients; the optimal rule can then be modeled nonparametrically from function and derivative evaluations within an RKHS.
2.3.1. Population and empirical problem
To allow a general situation as described above, consider a distribution $\mathbb{P}$ on $\mathcal{Z}\mathrel{\mathrel{\mathop{:}}=}\mathcal{X}\times\mathcal{Y}$, where $\mathcal{Y}\subset\mathbb{R}$. Let $s\in\mathbb{N}$. For any smooth function $h\in\mathcal{H}\subset\mathscr{C}^{s}(\mathcal{X})$, we define our target functional of interest (that depends on $h$ and its gradients up to order $s$) as:
(2.3) $\mathcal{R}(h;\boldsymbol{z})\mathrel{\mathrel{\mathop{:}}=}\sum_{\boldsymbol{\alpha}\in\mathcal{A}_{s}}w_{\boldsymbol{\alpha}}(\boldsymbol{z})D^{\boldsymbol{\alpha}}h(\boldsymbol{x})=\sum_{\boldsymbol{\alpha}\in\mathcal{A}_{s}}w_{\boldsymbol{\alpha}}(\boldsymbol{z})\langle h,\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x})\rangle_{\mathcal{H}}=\langle h,\psi(\boldsymbol{z})\rangle_{\mathcal{H}},$
where the weight coefficients $w_{\boldsymbol{\alpha}}(\boldsymbol{z})\in\mathbb{R}$ are known measurable functions of the data, and
(2.4) $\psi(\boldsymbol{z})\mathrel{\mathrel{\mathop{:}}=}\sum_{\boldsymbol{\alpha}\in\mathcal{A}_{s}}w_{\boldsymbol{\alpha}}(\boldsymbol{z})\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x})\in\mathcal{H}$
is the random vector in the RKHS that acts as the representer of the target functional $\mathcal{R}(\cdot\,;\boldsymbol{z})$.
Remark 2.2.
In many decision problems, the target object is a score that depends linearly on the level and on the gradient information of an unknown function $h$. A generic specification is given by (2.3), where $\boldsymbol{z}=(\boldsymbol{x},y)$ denotes observed data and the weights $w_{\boldsymbol{\alpha}}(\cdot)$ encode the task via the dependencies on $h$ and its gradients. In the case, the target functional only depends on the gradient values/specified derivatives only, we can consider $\mathcal{A}\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{\alpha}\in\mathcal{A}_{s}:w_{\boldsymbol{\alpha}}\equiv 0\}$. Such a form allows for flexibility and covers, for example, portfolio rules or control scores that use the function value as a signal and gradient or curvature components as sensitivity adjustments. Since each $D^{\boldsymbol{\alpha}}h(\boldsymbol{x})$ is a bounded linear functional in an RKHS with a sufficiently smooth kernel, $\mathcal{R}(h;\boldsymbol{z})$ remains linear in $h$ and admits a representer $\psi(\boldsymbol{z})$.
Now, we consider a mean-variance objective as follows:
(2.5) $\underset{h\in\mathcal{H}}{\operatorname{argmax}}\,\,\mathbb{E}_{\boldsymbol{z}\sim\mathbb{P}}[\mathcal{R}(h;\boldsymbol{z})]-\frac{1}{2}\mathbb{V}_{\boldsymbol{z}\sim\mathbb{P}}[\mathcal{R}(h;\boldsymbol{z})].$
Similar to learning problems in an RKHS, we set up the above as a (Tikhonov) regularized convex problem in $\mathcal{H}$ as follows. For $\lambda>0$:
(2.6) $h_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\underset{h\in\mathcal{H}}{\operatorname{argmin}}\,\,J_{\lambda}(h)\mathrel{\mathrel{\mathop{:}}=}-\mathbb{E}[\mathcal{R}(h;\boldsymbol{z})]+\frac{1}{2}\mathbb{V}[\mathcal{R}(h;\boldsymbol{z})]+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2},$
where $\mathbb{E}[\cdot]$ and $\mathbb{V}[\cdot]$ are taken with respect to the population distribution $\mathbb{P}$. Next, we define the empirical problem, given observations $\{\boldsymbol{z}_{i}\mathrel{\mathrel{\mathop{:}}=}(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}\sim\mathbb{P}$ as:
(2.7) $\psi_{i}\mathrel{\mathrel{\mathop{:}}=}\psi(\boldsymbol{z}_{i})=\sum_{\boldsymbol{\alpha}\in\mathcal{A}_{s}}w_{\boldsymbol{\alpha}}(\boldsymbol{z}_{i})\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x}_{i})\in\mathcal{H}.$
Then the empirical counterpart to Problem 2.6 is given by:
(2.8) $\widehat{h}_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\underset{h\in\mathcal{H}}{\operatorname{argmin}}\,\,\widehat{J}_{\lambda}(h)\mathrel{\mathrel{\mathop{:}}=}-\widehat{\mathbb{E}}[\mathcal{R}(h;\boldsymbol{z})]+\frac{1}{2}\widehat{\mathbb{V}}[\mathcal{R}(h;\boldsymbol{z})]+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2},$
where we use the notation $\widehat{}\mathbb{E}[\cdot]$ and $\widehat{\mathbb{V}}$$\cdot$] to refer to the mean and variance of the empirical distribution $\widehat{}\mathbb{P}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{N}\sum_{i=1}^{N}\delta_{\boldsymbol{z}_{i}}$.
2.3.2. Formulation in RKHS
Using the embedding (2.3) of the target functional $\mathcal{R}(\cdot;\cdot)$ in $\mathcal{H}$, we can formulate Problems 2.6 and 2.8 in the RKHS. We define the moments of the embedding with respect to the population and empirical distribution as follows:
[…]
We can compute the corresponding mean and variance of the target functional as follows:
(2.10) $\begin{split}\mathbb{E}[\mathcal{R}(h;\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{n})]&\mathrel{\mathrel{\mathop{:}}=}\mathbb{E}[\langle h,\psi_{i}\rangle_{\mathcal{H}}]=\langle h,\mathbb{E}[\psi_{i}]\rangle_{\mathcal{H}}=\langle h,\mu\rangle_{\mathcal{H}}\\ \mathbb{V}[\mathcal{R}(h;\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{n})]&\mathrel{\mathrel{\mathop{:}}=}\mathbb{E}[\langle h,\psi_{i}-\mu\rangle_{\mathcal{H}}^{2}]=\mathbb{E}[\langle h,\bigl((\psi_{i}-\mu)\otimes(\psi_{i}-\mu)\bigr)h\rangle_{\mathcal{H}}]=\langle h,\mathnormal{\Sigma}h\rangle_{\mathcal{H}}\\ \widehat{\mathbb{E}}[\mathcal{R}(h;\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{n})]&\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathbb{E}}[\langle h,\psi_{i}\rangle_{\mathcal{H}}]=\langle h,\widehat{\mathbb{E}}[\psi_{i}]\rangle_{\mathcal{H}}=\langle h,\widehat{\mu}\rangle_{\mathcal{H}}\\ \widehat{\mathbb{V}}[\mathcal{R}(h;\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{n})]&\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathbb{E}}[\langle h,\psi_{i}-\widehat{\mu}\rangle_{\mathcal{H}}^{2}]=\widehat{\mathbb{E}}[\langle h,(\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})h\rangle_{\mathcal{H}}]=\langle h,\widehat{\mathnormal{\Sigma}}h\rangle_{\mathcal{H}}.\end{split}$
Using the above characterizations, (2.3), we can write Problems 2.6 and 2.8 in terms of the quantities in (2.10). The equivalent representation of the population problem, cp. Problem 2.6 reads:
(2.11) $h_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\underset{h\in\mathcal{H}}{\operatorname{argmin}}\,\,J_{\lambda}(h)=-\langle h,\mu\rangle_{\mathcal{H}}+\frac{1}{2}\langle h,\mathnormal{\Sigma}h\rangle_{\mathcal{H}}+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2},$
whose empirical counterpart is
(2.12) $\widehat{h}_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\underset{h\in\mathcal{H}}{\operatorname{argmin}}\,\,\widehat{J}_{\lambda}(h)=-\langle h,\widehat{\mu}\rangle_{\mathcal{H}}+\frac{1}{2}\langle h,\widehat{\mathnormal{\Sigma}}h\rangle_{\mathcal{H}}+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2}.$
Remark 2.3.
The reproducing property of the derivatives of the feature function $\phi$ allows us to represent the target functional $\mathcal{R}(\cdot;\cdot)$ as the function $\psi$ in the RKHS $\mathcal{H}$. Such a nonparametric representation facilitates the characterization of the optimizers of Problems 2.6 and 2.8 in the RKHS $\mathcal{H}$, as above. We also remark that such a formulation also enables us to compute the closed-form expressions of the respective optimizers in terms of the quantities defined in (2.9). In particular, we can derive the representer theorem, see Theorem 2.4, for the empirical case, which leads to a finite system of equations, see (D.9).
2.4. Representer theorem
Since the goal is to find the optimal rule, the representer theorem helps us identify the specific subspace of $\mathcal{H}$ that contains it. Our version of the representer theorem, see Theorem 2.4 below, is a specialized case of (Zhou, 2008, Theorem 2).
Theorem 2.4 (Representer theorem).
The optimal solution to Problem 2.12 has the form
(2.13) $\widehat{h}_{\lambda}=\sum_{i=1}^{N}\sum_{\boldsymbol{\alpha}\in\mathcal{A}_{s}}\widehat{c}_{i,\boldsymbol{\alpha}}\,\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x}_{i}).$
Proof of Theorem 2.4.
By Theorem 2.1, for any $\boldsymbol{\alpha}\in\mathcal{A}_{s}$, $\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x})\in\mathcal{H}$. Define the subspace of $\mathcal{H}$ spanned by the feature function and its derivative evaluations at the sample points:
(2.14) $\mathcal{H}_{X}\mathrel{\mathrel{\mathop{:}}=}\operatorname{span}\Bigl\{\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x}_{i}):1\leq i\leq N,\,\boldsymbol{\alpha}\in\mathcal{A}_{s}\Bigr\}.$
$\mathcal{H}_{X}$ is a finite-dimensional closed subspace of $\mathcal{H}$, and therefore, we have the direct sum decomposition $\mathcal{H}=\mathcal{H}_{X}\oplus\mathcal{H}_{X}^{\perp}$. Hence, for any $h\in\mathcal{H}$, we can write $h=h_{0}+h_{1}$ with $h_{1}\perp\mathcal{H}_{X}$. Using (2.7), (2.9), and (2.14), $\widehat{\mu}\in\mathcal{H}_{X}$ and therefore $\psi_{i}-\widehat{\mu}\in\mathcal{H}_{X}$. This implies, from (2.9), that for $h\in\mathcal{H}$ with the above direct sum decomposition, the empirical covariance operator $\widehat{\mathnormal{\Sigma}}$ satisfies
$\widehat{\mathnormal{\Sigma}}(h_{0}+h_{1})=\widehat{\mathnormal{\Sigma}}h_{0}+\widehat{\mathnormal{\Sigma}}h_{1}=\widehat{\mathnormal{\Sigma}}h_{0}+\widehat{\mathbb{E}}[\langle h_{1},\psi_{i}-\widehat{\mu}\rangle_{\mathcal{H}}(\psi_{i}-\widehat{\mu})]=\widehat{\mathnormal{\Sigma}}h_{0}.$
This shows that for any $h\in\mathcal{H}$, the quadratic form defined by the empirical covariance operator depends on the orthogonal projection of $h$ onto the working subspace. Therefore, for any $h\in\mathcal{H}$,
$\widehat{J}_{\lambda}(h)$ $=-\langle h_{0}+h_{1},\psi_{i}\rangle_{\mathcal{H}}+\frac{1}{2}\langle h_{0}+h_{1},\widehat{\mathnormal{\Sigma}}(h_{0}+h_{1})\rangle_{\mathcal{H}}+\frac{\lambda}{2}\|h_{0}+h_{1}\|^{2}_{\mathcal{H}}$
$=-\langle h_{0},\psi_{i}\rangle_{\mathcal{H}}+\langle h_{0},\widehat{\mathnormal{\Sigma}}h_{0}\rangle_{\mathcal{H}}+\langle h_{1},\widehat{\mathnormal{\Sigma}}h_{0}\rangle_{\mathcal{H}}+\frac{\lambda}{2}\|h_{0}+h_{1}\|^{2}_{\mathcal{H}}$
$=-\langle h_{0},\psi_{i}\rangle_{\mathcal{H}}+\langle h_{0},\widehat{\mathnormal{\Sigma}}h_{0}\rangle_{\mathcal{H}}+\frac{\lambda}{2}\|h_{0}\|_{\mathcal{H}}^{2}+\frac{\lambda}{2}\|h_{1}\|^{2}_{\mathcal{H}}\quad\text{(by Pythagoras theorem)}$
$=\widehat{J}_{\lambda}(h_{0})+\frac{\lambda}{2}\|h_{1}\|^{2}_{\mathcal{H}}\geq\widehat{J}_{\lambda}(h_{0}).$
The above calculation shows that the value of the objective function for any $h\in\mathcal{H}$ is at least as large as its orthogonal projection onto $\mathcal{H}_{X}$. This leads to the expression as in (2.13). ∎
Remark 2.5.
As noted in Remark 2.2, when function values or certain derivatives are not used in (2.3), let $\mathcal{A}\subset\mathcal{A}_{s}$ denote the set of multi-indices that appear in the functional (equivalently, take $w_{\boldsymbol{\alpha}}\equiv 0$ for $\boldsymbol{\alpha}\notin\mathcal{A}$). In this case, the optimal solution is contained within the finite-dimensional space $\operatorname{span}\bigl\{\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x}_{i}):1\leq i\leq N,\,\boldsymbol{\alpha}\in\mathcal{A}\bigr\}$. The proof of Theorem 2.4 is valid without modification since $\psi_{i}$ and $\widehat{\mu}$ belong to this subspace, as does the minimizer $\widehat{h}_{\lambda}$. In particular, if there is no function-value term, then no $\phi(\boldsymbol{x}_{i})$ terms appear in the representer theorem.
The explicit form of $\widehat{h}_{\lambda}$ in (2.13) implies that the optimal solution to Problem 2.8 is parameterized by the optimal coefficients $\widehat{c}_{i,\boldsymbol{\alpha}}$. Thus, we need to find the optimal coefficients to evaluate the sample estimator at any point $\boldsymbol{x}\in\mathcal{X}$. This is done via solving for a finite system of equations, the details of which are deferred to Appendix D.
3. Statistical properties of sample estimator
This section develops the statistical properties of the estimator $\widehat{h}_{\lambda}$ from (2.12). We begin by stating the assumptions, followed by establishing the asymptotic properties and the finite-sample deviation bounds, in Propositions 3.4 and Proposition 3.5, respectively.
3.1. Assumptions and setting
We begin with the assumption that $\psi_{i}\in\mathcal{H}$ from (2.7) are independently and identically distributed (i.i.d). Note that if the observations $\{\boldsymbol{z}_{i}\mathrel{\mathrel{\mathop{:}}=}(\boldsymbol{x}_{i},y_{i})\}_{i=1}^{N}$ are i.i.d., then due to measurability of the weight functions $w_{\boldsymbol{\alpha}}(\cdot)$, the feature function and its derivatives $\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x}_{i})$, the random vectors $\psi_{i}\in\mathcal{H}$ will be i.i.d. as well. Next, we define the Hilbert space
(3.1) $\mathbb{H}\mathrel{\mathrel{\mathop{:}}=}\mathcal{H}\oplus\mathscr{H}\mathscr{S}(\mathcal{H}),$
equipped with the inner product
[…]
Since $\mathcal{H}$ is separable, the space of Hilbert-Schmidt operators $\mathscr{H}\mathscr{S}(\mathcal{H})$ is also separable, see Bosq (2000, Chapter 1). This implies that $\mathbb{H}$ is also a separable Hilbert space. We denote by $\widetilde{\psi}_{i}\mathrel{\mathrel{\mathop{:}}=}\psi_{i}-\mu$, the centered $\mathcal{H}$-valued random vectors, i.e., $\mathbb{E}[\widetilde{\psi}_{i}]=0$, and we define the covariance operators:
(3.3) $\widetilde{\mathnormal{\Sigma}}\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathbb{E}}[\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}]=\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i},\qquad\widetilde{\mathcal{C}}_{i}\mathrel{\mathrel{\mathop{:}}=}\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}-\mathnormal{\Sigma}.$
We have the following proposition, which we prove in Appendix A.
Proposition 3.1 (Moment and operator properties).
Let $\psi_{i}$ defined in (2.7) be i.i.d. in the separable Hilbert space $\mathcal{H}$. Let $\widetilde{\psi}_{i}\mathrel{\mathrel{\mathop{:}}=}\psi_{i}-\mu,\,\mathnormal{\Sigma}=\mathbb{E}[\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}]$, and $\widetilde{\mathnormal{\Sigma}},\,\widetilde{}\mathcal{C}_{i}$ be defined as in (3.3). Under the assumption $\mathbb{E}\|\psi_{1}\|_{\mathcal{H}}^{4}<\infty$, the following hold,
(3.4) $\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}<\infty,\qquad\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}<\infty.$
Moreover, the covariance operators $\mathnormal{\Sigma},\,\widehat{\mathnormal{\Sigma}},\,\widetilde{\mathnormal{\Sigma}}$ are positive, self-adjoint, and Hilbert-Schmidt, while each $\widetilde{}\mathcal{C}_{i}$ is self-adjoint and Hilbert-Schmidt with $\mathbb{E}[\widetilde{\mathcal{C}}_{i}]=0$.
Next, we have the following identities for the zero-mean processes $\widetilde{\psi}_{i},\,\widetilde{}\mathcal{C}_{i}$.
(3.5) $\begin{split}\widehat{\mu}-\mu&=\frac{1}{N}\sum_{i=1}^{N}\psi_{i}-\mu=\frac{1}{N}\sum_{i=1}^{N}(\psi_{i}-\mu)=\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}_{i},\\ \widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma}&=\frac{1}{N}\sum_{i=1}^{N}\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}-\mathnormal{\Sigma}=\frac{1}{N}\sum_{i=1}^{N}(\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}-\mathnormal{\Sigma})=\frac{1}{N}\sum_{i=1}^{N}\widetilde{\mathcal{C}}_{i}.\end{split}$
We shall henceforth use the notation $\mathnormal{\Sigma}_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\mathnormal{\Sigma}+\lambda I,\;\widehat{\mathnormal{\Sigma}}_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathnormal{\Sigma}}+\lambda I$ for the remainder of the paper. Since $\lambda>0$, the eigenvalues of $\mathnormal{\Sigma}_{\lambda}$ and $\widehat{\mathnormal{\Sigma}}_{\lambda}$ are bounded away from zero and the operators $\mathnormal{\Sigma}_{\lambda}^{-1},\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\in\mathscr{B}(\mathcal{H})$.
3.2. Asymptotic properties
In order to show the asymptotic results, we start by characterizing the optimal solutions to the Problems 2.11 and 2.12.
Proposition 3.2 (Optimal solution).
Let $h_{\lambda}$ and $\widehat{h}_{\lambda}$ be defined as the optimal solution to Problem 2.11 and Problem 2.12 respectively. Then, it holds,
(3.6) $h_{\lambda}=\mathnormal{\Sigma}_{\lambda}^{-1}\mu,\qquad\widehat{h}_{\lambda}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\widehat{\mu}$.
Now, from Lemma A.2, we have the following decomposition:
(3.7) $\widehat{h}_{\lambda}-h_{\lambda}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}((\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda})+r_{N},\qquad r_{N}\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}(\widetilde{\mathnormal{\Sigma}}-\widehat{\mathnormal{\Sigma}})h_{\lambda}.$
Equation 3.7 shows that the error decomposes as the sum of a main fluctuation and a remainder term. In what follows, we seek to show that the fluctuation term is asymptotically gaussian, see Proposition 3.4, while the remainder term decays as $o_{\mathbb{P}}(N^{-1/2})$, see Lemma A.4. To show the former, we first use a functional central limit theorem (CLT) applicable for i.i.d. sequences in separable Hilbert spaces, and then use the continuous mapping theorem (CMT). Towards that end, we define the following function on $\mathbb{H}$,
(3.8) $F\colon\mathbb{H}\to\mathcal{H},\qquad F(h,\mathcal{C})\mathrel{\mathrel{\mathop{:}}=}h-\mathcal{C}h_{\lambda}\quad\text{for }(h,\mathcal{C})\in\mathbb{H}.$
$F$ is a bounded linear map on $\mathcal{H}$, see Lemma A.5. We now state and prove the following.
Proposition 3.3 (Asymptotic gaussianity).
Under the assumptions of Proposition 3.1, it holds,
(3.9) $\sqrt{N}\Bigl((\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda}\Bigr)\overset{d}{\longrightarrow}\mathcal{N}_{\mathcal{H}}(0,\mathcal{Q}_{\lambda}),$
where
(3.10) $\mathcal{Q}_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\mathbb{E}[F(\widetilde{\psi}_{1},\widetilde{\mathcal{C}}_{1})\otimes F(\widetilde{\psi}_{1},\widetilde{\mathcal{C}}_{1})]=\mathbb{E}[(\widetilde{\psi}_{1}-\widetilde{\mathcal{C}}_{1}h_{\lambda})\otimes(\widetilde{\psi}_{1}-\widetilde{\mathcal{C}}_{1}h_{\lambda})].$
Proof.
From Proposition 3.1, $\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}<\infty$. Denote the sample mean as
(3.11) $\bar{S}_{N}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{N}\sum_{i=1}^{N}(\widetilde{\psi}_{i},\widetilde{\mathcal{C}}_{i})=(\widehat{\mu}-\mu,\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma}),$
where the last equality follows from (3.5). From Bosq (2000, Theorem 2.7), the CLT holds,
(3.12) $\sqrt{N}\bar{S}_{N}=\sqrt{N}(\widehat{\mu}-\mu,\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})\overset{d}{\longrightarrow}\mathcal{N}_{\mathbb{H}}(0,\Gamma),\qquad\Gamma\mathrel{\mathrel{\mathop{:}}=}\mathbb{E}[(\widetilde{\psi}_{1},\widetilde{\mathcal{C}}_{1})\otimes(\widetilde{\psi}_{1},\widetilde{\mathcal{C}}_{1})].$
We now use $F$ from (3.8) to apply the CMT to (3.12) ,
$\sqrt{N}\Bigl((\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda}\Bigr)\overset{d}{\longrightarrow}\mathcal{N}_{\mathcal{H}}\Bigl(0,\mathcal{Q}_{\lambda}\Bigr),$
where $\mathcal{Q}_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\mathbb{E}[F(\widetilde{\psi}_{1},\widetilde{\mathcal{C}}_{1})\otimes F(\widetilde{\psi}_{1},\widetilde{\mathcal{C}}_{1})]=\mathbb{E}[(\widetilde{\psi}_{1}-\widetilde{\mathcal{C}}_{1}h_{\lambda})\otimes(\widetilde{\psi}_{1}-\widetilde{\mathcal{C}}_{1}h_{\lambda})]$. ∎
Proposition 3.3 facilitates us to derive the asymptotic distribution of the main fluctuation term in (3.7). We are now ready to state the following lemma.
Proposition 3.4 (Asymptotic properties).
Under the assumptions of Proposition 3.1, it holds,
$(i)\;\widehat{h}_{\lambda}\overset{a.s.}{\longrightarrow}h_{\lambda}\qquad(ii)\;\sqrt{N}\left(\widehat{h}_{\lambda}-h_{\lambda}\right)\overset{d}{\longrightarrow}\mathcal{N}_{\mathcal{H}}(0,\mathcal{C}_{\lambda}),\qquad\mathcal{C}_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\mathnormal{\Sigma}_{\lambda}^{-1}\mathcal{Q}_{\lambda}\mathnormal{\Sigma}_{\lambda}^{-1}.$
Proof.
$(i)$ We have $\|\widehat{\mu}-\mu\|_{\mathcal{H}}\overset{a.s.}{\longrightarrow}0$ and $\|\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\text{op}}\leq\|\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}0$ from Lemma A.1. The map $A\mapsto(A+\lambda I)^{-1}$ is continuous in the operator norm topology on $\mathscr{B}(\mathcal{H})$ and hence, by CMT,
(3.13) $\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\overset{a.s.}{\longrightarrow}\mathnormal{\Sigma}_{\lambda}^{-1}.$
By the continuity of the bilinear map $(A,v)\mapsto Av$ on $\mathscr{B}(\mathcal{H})\times\mathcal{H}$, it follows from CMT that
$\widehat{h}_{\lambda}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\widehat{\mu}\overset{a.s.}{\longrightarrow}\mathnormal{\Sigma}_{\lambda}^{-1}\mu=h_{\lambda}.$
$(ii)$ $\sqrt{N}\,r_{N}\overset{\mathbb{P}}{\longrightarrow}0$ from Lemma A.4, while $\sqrt{N}((\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda})\overset{d}{\longrightarrow}\mathcal{N}_{\mathcal{H}}(0,\mathcal{Q}_{\lambda})$ from (3.9). Again, $\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\overset{\mathbb{P}}{\longrightarrow}\mathnormal{\Sigma}_{\lambda}^{-1}$ follows from the proof of asymptotic consistency above. Hence, by Slutsky’s theorem,
(3.14) $\begin{split}\sqrt{N}\left(\widehat{h}_{\lambda}-h_{\lambda}\right)&=\underbrace{\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}}_{\overset{\mathbb{P}}{\longrightarrow}\mathnormal{\Sigma}_{\lambda}^{-1}}\cdot\underbrace{\sqrt{N}\left((\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda}\right)}_{\overset{d}{\longrightarrow}\mathcal{N}_{\mathcal{H}}(0,\mathcal{Q}_{\lambda})}+\underbrace{\sqrt{N}\,r_{N}}_{\overset{\mathbb{P}}{\longrightarrow}0}\\ &\overset{d}{\longrightarrow}\mathcal{N}_{\mathcal{H}}(0,\mathnormal{\Sigma}_{\lambda}^{-1}\mathcal{Q}_{\lambda}\mathnormal{\Sigma}_{\lambda}^{-1}).\end{split}$
3.3. Finite-sample properties
We state the main result of this section below.
Proposition 3.5 (Finite-sample deviation bound).
Under the assumptions of Proposition 3.1, it holds with sampling probability at least $(1-\delta)$,
$\|\widehat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}\leq C_{FS}\left(\delta,\|h_{\lambda}\|_{\mathcal{H}}\right)\lambda^{-1}N^{-1/2},$
for the coefficient where
(3.15) $C(\delta,s)\mathrel{\mathrel{\mathop{:}}=}\sqrt{1+s^{2}}\,\sqrt{\frac{2\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{\mathcal{C}}_{1})\|_{\mathbb{H}}^{2}}{\delta}}+2s\,\frac{\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}}{\delta}.$
Proof.
Since $(\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda}=F(\bar{S}_{N})$, hence, (3.7) implies
$\|\widehat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}\leq\|\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}((\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda})\|_{\mathcal{H}}+\|r_{N}\|_{\mathcal{H}}=\|\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}F(\bar{S}_{N})\|_{\mathcal{H}}+\|r_{N}\|_{\mathcal{H}}.$
Now, $\|\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\|_{\text{op}}\leq 1/\lambda$, see the proof of Lemma A.4. Moreover, $\|F\|_{\text{op}}\leq\sqrt{1+\|h_{\lambda}\|_{\mathcal{H}}^{2}}$ from Lemma A.5 and $\|r_{N}\|_{\mathcal{H}}\leq\|h_{\lambda}\|_{\mathcal{H}}\,\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}/\lambda$ from (A.6). Hence, we have
$\|\widehat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}\leq\|\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\|_{\text{op}}\,\|F\|_{\text{op}}\,\|\bar{S}_{N}\|_{\mathbb{H}}+\|r_{N}\|_{\mathcal{H}}\leq\frac{\sqrt{1+\|h_{\lambda}\|_{\mathcal{H}}^{2}}}{\lambda}\|\bar{S}_{N}\|_{\mathbb{H}}+\frac{\|h_{\lambda}\|_{\mathcal{H}}}{\lambda}\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}.$
Choose $\delta\in(0,1)$. Then, with probability at most $\delta/2$, it holds, $\|\bar{S}_{N}\|_{\mathbb{H}}>\sqrt{2\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}/\delta}$, see Lemma A.6 and $\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}>2\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}/N\delta$, see Lemma A.3. Therefore, combining the probabilities via a union bound gives that with probability at least $(1-\delta)$,
$\|\widehat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}$ $\leq\frac{\sqrt{1+\|h_{\lambda}\|_{\mathcal{H}}^{2}}}{\lambda}\,\sqrt{\frac{2\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}}{\delta}}+\frac{\|h_{\lambda}\|_{\mathcal{H}}}{\lambda}\frac{2\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}}{N\delta}$
$\leq\frac{1}{\lambda\sqrt{N}}\underbrace{\left(\sqrt{1+\|h_{\lambda}\|_{\mathcal{H}}^{2}}\,\sqrt{\frac{2\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}}{\delta}}+2\|h_{\lambda}\|_{\mathcal{H}}\,\frac{\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}}{\delta}\right)}_{\mathrel{\mathrel{\mathop{:}}=}C_{FS}(\delta,\|h_{\lambda}\|_{\mathcal{H}})}$
$=C_{FS}(\delta,\|h_{\lambda}\|_{\mathcal{H}})\lambda^{-1}N^{-1/2}.$
Remark 3.6.
Proposition 3.5 shows that the estimation error $\|\widehat{h}_{\lambda}-h_{\lambda}\|$ admists a high-probability control of order $\mathcal{O}_{\mathbb{P}}(\lambda^{-1}N^{-1/2})$. The explicit constant $C_{FS}$ depends on the level $\delta\in(0,1)$, the size of the population solution $\|h_{\lambda}\|_{\mathcal{H}}$, the variance $\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}$, and the joint variance $\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}$. In particular, the rate matches the classical Monte Carlo $N^{-1/2}$ rate, with an additional $\lambda^{-1}$ term reflecting regularization. Thus, even without strong boundedness/tail assumptions (only a finite fourth moment assumption is required), we obtain non-asymptotic guarantees that complement the asymptotic results in Proposition 3.4.
4. Statistical Inference for shape constraints
In this section, we describe statistical inference for shape constraints of the sample estimator $\widehat{h}_{\lambda}$, with a focus on directional tests. Although our framework admits full multi-index differentiation on $\mathcal{X}\subset\mathbb{R}^{d}$, in many applications it is natural to assess shape restrictions along a fixed coordinate direction in the covariate space; for example, classic constraints in a fixed coordinate direction with small $s$: positivity corresponds to order 0, monotonicity to first order, and convexity to second order in the chosen direction. However, for an easier reading, we keep the general structure unchanged and discuss directional tests in Section 5.
4.1. Test statistic
We start by choosing any derivative order $\boldsymbol{\alpha}\in\mathcal{A}_{s}$, followed by deriving the necessary asymptotic results which help us arrive at the asymptotic distribution of the test statistic. The construction of the test statistic proceeds in the same way for any $\boldsymbol{\alpha}\in\mathcal{A}_{s}$. Hence, in what follows, we define the relevant quantities without attributing to the derivative order.
4.1.1. Setting
Recall that the sample estimator $\widehat{h}_{\lambda}\in\mathcal{H}\subset\mathscr{C}^{s}(\mathcal{X})$ for some fixed $s\in\mathbb{N}$. Choose $\boldsymbol{\alpha}\in\mathcal{A}_{s}$. For any finite testing grid $\mathcal{G}=\{\boldsymbol{\xi}_{j}\}_{j=1}^{n}\subset\mathcal{X}$, consider the vector of evaluations:
(4.1) $\boldsymbol{\theta}\mathrel{\mathrel{\mathop{:}}=}\big[h_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\big]_{j=1}^{n}\in\mathbb{R}^{n},\qquad\widehat{\boldsymbol{\theta}}\mathrel{\mathrel{\mathop{:}}=}\big[\widehat{h}_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\big]_{j=1}^{n}\in\mathbb{R}^{n}.$
From the reproducing property of the derivatives, see Theorem 2.1, it follows that for each $\boldsymbol{\xi}_{j}$, the evaluation of the partial derivatives may be represented via $\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})$ as $h^{\boldsymbol{\alpha}}(\boldsymbol{\xi}_{j})=\langle h,\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}$ for any $h\in\mathcal{H}$. We define the corresponding population and sample quantities:
(4.2) $u_{j}\mathrel{\mathrel{\mathop{:}}=}\mathnormal{\Sigma}_{\lambda}^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j}),\qquad\widehat{u}_{j}\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j}),\quad 1\leq j\leq n.$
Now, we define the following:
(4.3) $\widehat{F}_{i}\mathrel{\mathrel{\mathop{:}}=}(\psi_{i}-\widehat{\mu})-\bigl((\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-\widehat{\mathnormal{\Sigma}}\bigr)\widehat{h}_{\lambda},\qquad\widehat{}\mathcal{Q}_{\lambda}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{N}\sum_{i=1}^{N}\widehat{F}_{i}\otimes\widehat{F}_{i}.$
Note from (3.9) that the population analogue of $\widehat{}\mathcal{Q}_{\lambda}$ is $\mathcal{Q}_{\lambda}=\mathbb{E}[F_{i}\otimes F_{i}]$, where
(4.4) $F_{i}\mathrel{\mathrel{\mathop{:}}=}F(\widetilde{\psi}_{i},\widetilde{}\mathcal{C}_{i})=\widetilde{\psi}_{i}-\widetilde{}\mathcal{C}_{i}h_{\lambda}=(\psi_{i}-\mu)-\bigl((\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\bigr)h_{\lambda}.$
With these quantities at hand, we define the $n\times n$ covariance matrices with pairwise entries
(4.5) $\begin{split}[\boldsymbol{\Omega}_{\lambda}]_{k,j}&\mathrel{\mathrel{\mathop{:}}=}\langle u_{k},\mathcal{Q}_{\lambda}u_{j}\rangle_{\mathcal{H}}=\mathbb{E}\left[\langle F_{i},u_{k}\rangle_{\mathcal{H}}\,\langle F_{i},u_{j}\rangle_{\mathcal{H}}\right],\\ [\widehat{\boldsymbol{\Omega}}_{\lambda}]_{k,j}&\mathrel{\mathrel{\mathop{:}}=}\langle\widehat{u}_{k},\widehat{\mathcal{Q}}_{\lambda}\widehat{u}_{j}\rangle_{\mathcal{H}}=\frac{1}{N}\sum_{i=1}^{N}\langle\widehat{F}_{i},\widehat{u}_{k}\rangle_{\mathcal{H}}\,\langle\widehat{F}_{i},\widehat{u}_{j}\rangle_{\mathcal{H}}.\end{split}$
Finally, we define the bounded linear operator that evaluates the derivative at the grid points,
(4.6) $\mathcal{S}_{n}\colon\mathcal{H}\to\mathbb{R}^{n},\qquad\mathcal{S}_{n}(h)\mathrel{\mathrel{\mathop{:}}=}\Bigl[\langle h,\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}\Bigr]_{j=1}^{n}\in\mathbb{R}^{n},$
whose adjoint is given as
(4.7) $\mathcal{S}_{n}^{\ast}\colon\mathbb{R}^{n}\to\mathcal{H},\qquad\mathcal{S}_{n}^{\ast}(\boldsymbol{\omega})\mathrel{\mathrel{\mathop{:}}=}\sum_{j=1}^{n}\omega_{j}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j}).$
4.1.2. Asymptotic properties
We first establish the large-sample behavior of the derivative evaluations on the grid. This result underpins inference for shape constraints.
Proposition 4.1 (Asymptotic distribution).
Let $\boldsymbol{\theta},\,\widehat{\boldsymbol{\theta}}$ be defined as in (4.1). Under the assumptions of Proposition 3.1, it holds,
(4.8) $\sqrt{N}\Bigl(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}\Bigr)\overset{d}{\longrightarrow}\mathcal{N}_{n}(\boldsymbol{0},\boldsymbol{\Omega}_{\lambda}).$
Proof.
By (4.6) and Theorem 2.1, we have
$\mathcal{S}_{n}(\widehat{h}_{\lambda}-h_{\lambda})=\Bigl[\langle\widehat{h}_{\lambda}-h_{\lambda},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}\Bigr]_{j=1}^{n}=\Bigl[\widehat{h}_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})-h_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\Bigr]_{j=1}^{n}\in\mathbb{R}^{n}.$
From (ii) of Proposition 3.4, we have $\sqrt{N}\Bigl(\widehat{h}_{\lambda}-h_{\lambda}\Bigr)\overset{d}{\longrightarrow}\mathcal{N}_{\mathcal{H}}(0,\mathcal{C}_{\lambda})$ where $\mathcal{C}_{\lambda}=\mathnormal{\Sigma}_{\lambda}^{-1}\mathcal{Q}_{\lambda}\mathnormal{\Sigma}_{\lambda}^{-1}$. Using CMT, we obtain
$\sqrt{N}\Bigl(\widehat{\boldsymbol{\theta}}-\boldsymbol{\theta}\Bigr)=\sqrt{N}\Bigl[\widehat{h}_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})-h_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\Bigr]_{j=1}^{n}=\sqrt{N}\,\mathcal{S}_{n}(\widehat{h}_{\lambda}-h_{\lambda})\overset{d}{\longrightarrow}\mathcal{N}_{n}(\boldsymbol{0},\mathcal{S}_{n}\mathcal{C}_{\lambda}\mathcal{S}_{n}^{\ast}).$
It remains to show that $\mathcal{S}_{n}\mathcal{C}_{\lambda}\mathcal{S}_{n}^{\ast}=\boldsymbol{\Omega}_{\lambda}$. First note that $\mathcal{S}_{n}\mathcal{C}_{\lambda}\mathcal{S}_{n}^{\ast}\colon\mathbb{R}^{n}\to\mathbb{R}^{n}$ is a bounded linear operator, hence we can represent its action as an $n\times n$ matrix with respect to the canonical basis of $\mathbb{R}^{n}$. Let the (canonical) basis vectors be denoted as $\{\boldsymbol{e}_{j}\}_{j=1}^{n}\in\mathbb{R}^{n}$. Then, for $1\leq k,\,j\leq n$,
$[\mathcal{S}_{n}\mathcal{C}_{\lambda}\mathcal{S}_{n}^{\ast}]_{k,j}=\langle\boldsymbol{e}_{k},\mathcal{S}_{n}\mathcal{C}_{\lambda}\mathcal{S}_{n}^{\ast}\,\boldsymbol{e}_{j}\rangle_{\mathbb{R}^{n}}=\langle\mathcal{S}_{n}^{\ast}\,\boldsymbol{e}_{k},\mathcal{C}_{\lambda}\mathcal{S}_{n}^{\ast}\,\boldsymbol{e}_{j}\rangle_{\mathcal{H}}=\langle\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{k}),\mathcal{C}_{\lambda}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}.$
Using the definition of $\mathcal{C}_{\lambda}$ (see Proposition 3.4) and $u_{j}$ from (4.2), $\mathcal{S}_{n}\mathcal{C}_{\lambda}\mathcal{S}_{n}^{\ast}]_{k,j}=\langle\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{k}),\mathnormal{\Sigma}_{\lambda}^{-1}\mathcal{Q}_{\lambda}\mathnormal{\Sigma}_{\lambda}^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}=\langle\mathnormal{\Sigma}_{\lambda}^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{k}),\mathcal{Q}_{\lambda}\mathnormal{\Sigma}_{\lambda}^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}=\langle u_{k},\mathcal{Q}_{\lambda}u_{j}\rangle_{\mathcal{H}}$, which proves the claim that $\mathcal{S}_{n}\mathcal{C}_{\lambda}\mathcal{S}_{n}^{\ast}=\boldsymbol{\Omega}_{\lambda}$. ∎
To make Proposition 4.1 feasible in practice, we need a consistent estimator of the asymptotic covariance matrix $\boldsymbol{\Omega}_{\lambda}$. The following result shows that the plug-in estimator $\widehat{\boldsymbol{\Omega}}_{\lambda}$ converges to its population analogue.
Theorem 4.2 (Consistency of covariance estimator).
Let $\boldsymbol{\Omega}_{\lambda},\,\widehat{\boldsymbol{\Omega}}_{\lambda}$ be defined as in (4.5). Under the assumptions of Proposition 3.1, it holds,
(4.9) $\widehat{\boldsymbol{\Omega}}_{\lambda}\overset{a.s.}{\longrightarrow}\boldsymbol{\Omega}_{\lambda}\quad\text{as}\quad N\to\infty.$
Proof.
For any $1\leq k,j\leq n$, we have $\left|[\widehat{\boldsymbol{\Omega}}_{\lambda}]_{k,j}-[\boldsymbol{\Omega}_{\lambda}]_{k,j}\right|=\Big|\langle\widehat{u}_{k},\widehat{\mathcal{Q}}_{\lambda}\widehat{u}_{j}\rangle_{\mathcal{H}}-\langle u_{k},\mathcal{Q}_{\lambda}u_{j}\rangle_{\mathcal{H}}\Big|$. We can decompose the error as $\langle\widehat{u}_{k},\widehat{\mathcal{Q}}_{\lambda}\widehat{u}_{j}\rangle_{\mathcal{H}}-\langle u_{k},\mathcal{Q}_{\lambda}u_{j}\rangle_{\mathcal{H}}=\langle\widehat{u}_{k}-u_{k},\widehat{\mathcal{Q}}_{\lambda}\widehat{u}_{j}\rangle_{\mathcal{H}}+\langle u_{k},(\widehat{\mathcal{Q}}_{\lambda}-\mathcal{Q}_{\lambda})\widehat{u}_{j}\rangle_{\mathcal{H}}+\langle u_{k},\mathcal{Q}_{\lambda}(\widehat{u}_{j}-u_{j})\rangle_{\mathcal{H}}$. Hence, using the triangle inequality,
$\left|[\widehat{\boldsymbol{\Omega}}_{\lambda}]_{k,j}-[\boldsymbol{\Omega}_{\lambda}]_{k,j}\right|\leq\underbrace{\left|\langle\widehat{u}_{k}-u_{k},\widehat{\mathcal{Q}}_{\lambda}\widehat{u}_{j}\rangle_{\mathcal{H}}\right|}_{(I)}+\underbrace{\left|\langle u_{k},(\widehat{\mathcal{Q}}_{\lambda}-\mathcal{Q}_{\lambda})\widehat{u}_{j}\rangle_{\mathcal{H}}\right|}_{(II)}+\underbrace{\left|\langle u_{k},\mathcal{Q}_{\lambda}(\widehat{u}_{j}-u_{j})\rangle_{\mathcal{H}}\right|}_{(III)}.$
Now, using (3.13) and the definition of $\widehat{u}_{j},\,u_{j}$ from (4.2), we have for any $j=1,\ldots,n$,
(4.10) $\|\widehat{u}_{j}-u_{j}\|_{\mathcal{H}}=\|\big(\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}-\mathnormal{\Sigma}_{\lambda}^{-1}\big)\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\|_{\mathcal{H}}\leq\underbrace{\|\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}-\mathnormal{\Sigma}_{\lambda}^{-1}\|_{\text{op}}}_{\overset{a.s.}{\longrightarrow}0}\;\underbrace{\|\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\|_{\mathcal{H}}}_{<\infty}\overset{a.s.}{\longrightarrow}0.$
From Lemma B.3, $\|\widehat{\mathcal{Q}}_{\lambda}-\mathcal{Q}_{\lambda}\|_{\text{op}}\leq\|\widehat{\mathcal{Q}}_{\lambda}-\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}0$, while Lemma B.2 gives $\|u_{j}\|_{\mathcal{H}},\,\|\widehat{u}_{j}\|_{\mathcal{H}},\,\|\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}<\infty$, and $\|\widehat{}\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}=\mathcal{O}(1)$ a.s. $N\to\infty$. Hence,
$\begin{split}(I)&\leq\|\widehat{u}_{k}-u_{k}\|_{\mathcal{H}}\,\|\widehat{}\mathcal{Q}_{\lambda}\|_{\text{op}}\,\|\widehat{u}_{j}\|_{\mathcal{H}}\overset{a.s.}{\longrightarrow}0,\\ (II)&\leq\|u_{k}\|_{\mathcal{H}}\,\|\widehat{\mathcal{Q}}_{\lambda}-\mathcal{Q}_{\lambda}\|_{\text{op}}\,\|\widehat{u}_{j}\|_{\mathcal{H}}\overset{a.s.}{\longrightarrow}0,\\ (III)&\leq\|u_{k}\|_{\mathcal{H}}\,\|\mathcal{Q}_{\lambda}\|_{\text{op}}\,\|\widehat{u}_{j}-u_{j}\|_{\mathcal{H}}\overset{a.s.}{\longrightarrow}0.\end{split}$
Therefore, for any $1\leq k,j\leq n$, it holds, $\widehat{\boldsymbol{\Omega}}_{\lambda}]_{k,j}\overset{a.s.}{\longrightarrow}[\boldsymbol{\Omega}_{\lambda}]_{k,j}$. Since the grid size $n$ is fixed, hence, $\widehat{\boldsymbol{\Omega}}_{\lambda}\overset{a.s.}{\longrightarrow}\boldsymbol{\Omega}_{\lambda}$ as $N\to\infty$. ∎
4.1.3. Test statistic
Theorem 4.2 shows the consistency of the finite-sample $n\times n$ covariance matrix $\widehat{\boldsymbol{\Omega}}_{\lambda}$. As a final step towards constructing the test statistic, we define the following. We refer an interested reader to Silvapulle and Sen (2001) for further details.
Definition 4.3 (Chi-bar-squared distribution).
Let $\mathcal{M}\subset\mathbb{R}^{n}$ be a closed convex cone, and let $\boldsymbol{Z}\sim\mathcal{N}_{n}(\boldsymbol{0},\boldsymbol{V})$ where $\boldsymbol{V}$ is a symmetric and positive definite matrix. Then, $\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})$ is defined to be the random variable having the same distribution as
(4.11) $\boldsymbol{Z}^{\top}\boldsymbol{V}^{-1}\boldsymbol{Z}-\underset{\boldsymbol{x}\in\mathcal{M}}{\min}\,\,(\boldsymbol{Z}-\boldsymbol{x})^{\top}\boldsymbol{V}^{-1}(\boldsymbol{Z}-\boldsymbol{x}).$
Denote by $\mathcal{M}^{\circ}\mathrel{\mathrel{\mathop{:}}=}\{\boldsymbol{x}:\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\boldsymbol{V}^{-1}}\leq 0\text{ for all $\boldsymbol{y}\in\mathcal{M}$}\}$ the polar cone of $\mathcal{M}$, where we define the inner product $\langle\boldsymbol{x},\boldsymbol{y}\rangle_{\boldsymbol{V}^{-1}}\mathrel{\mathrel{\mathop{:}}=}\boldsymbol{x}^{\top}\boldsymbol{V}^{-1}\boldsymbol{y}$. From Silvapulle and Sen (2001, Proposition 3.4.1), we have:
(4.12) $\|\Pi^{\boldsymbol{V}^{-1}}_{\mathcal{M}}(\boldsymbol{Z})\|_{\boldsymbol{V}^{-1}}^{2}\sim\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M}),\qquad\|\boldsymbol{Z}-\Pi^{\boldsymbol{V}^{-1}}_{\mathcal{M}}(\boldsymbol{Z})\|_{\boldsymbol{V}^{-1}}^{2}\sim\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M}^{\circ}),$
where $\Pi^{\boldsymbol{V}^{-1}}_{\mathcal{M}}(\boldsymbol{Z})$ is the orthogonal projection of $\boldsymbol{Z}$ onto $\mathcal{M}$ under the inner product $\langle\cdot,\cdot\rangle_{\boldsymbol{V}^{-1}}$. Moreover, we have:
Theorem 4.4 (Silvapulle and Sen (2001, Theorem 3.4.2)).
Let $\mathcal{M}$ be a closed convex cone in $\mathbb{R}^{n}$ and let $\boldsymbol{V}\in\mathbb{R}^{n\times n}$ be a symmetric and positive definite matrix. Then the distribution of $\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})$ is given by
(4.13) $\mathbb{P}\left(\bar{\chi}^{2}(\boldsymbol{V},\mathcal{M})\leq c\right)=\sum_{j=0}^{n}w_{j}(n,\boldsymbol{V},\mathcal{M})\,\mathbb{P}(\chi^{2}_{j}\leq c),$
where $w_{j}(n,\boldsymbol{V},\mathcal{M})\geq 0$ for $0\leq j\leq n$ and $\sum_{j=0}^{n}w_{j}(n,\boldsymbol{V},\mathcal{M})=1$.
We also put the following result, which characterizes an orthogonal projection in Hilbert spaces.
Theorem 4.5 (Bauschke and Combettes (2017, Theorem 3.16)).
Let $\mathcal{M}$ be a non-empty closed convex subset of a Hilbert space $\mathscr{H}$. Then for any $u\in\mathscr{H}$, the orthogonal projection $\Pi_{\mathcal{M}}(u)$ (under the $\mathscr{H}$-inner product) is well-defined and unique and satisfies
(4.14) $\langle u-\Pi_{\mathcal{M}}(u),v-\Pi_{\mathcal{M}}(u)\rangle_{\mathscr{H}}\leq 0\quad\text{for any }v\in\mathcal{M}.$
For a fixed derivative order $\boldsymbol{\alpha}\in\mathcal{A}_{s}$ and a grid $\mathcal{G}\subset\mathcal{X}$, we test the one-sided composite cone restriction given by the positivity constraint of the $\boldsymbol{\alpha}$-derivative evaluation at the grid points:
(4.15) $H_{0}:\boldsymbol{\theta}=[h_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})]_{j=1}^{n}\in\mathbb{R}^{n}_{+}\;\text{vs.}\;H_{1}:\,\text{there exists some $j\in\{1,\ldots,n\}$ such that $h_{\lambda}^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})<0$}.$
The least favorable null is given by $\boldsymbol{\theta}=\boldsymbol{0}$, that is, the boundary of the positive orthant $\mathbb{R}^{n}_{+}$. We now state the following theorem, which defines the test statistic and shows its asymptotic distribution, see Appendix C for a proof.
Theorem 4.6 (Test statistic).
Define the test statistic
(4.16) $W_{N}\mathrel{\mathrel{\mathop{:}}=}\underset{\boldsymbol{c}\in\mathbb{R}^{n}_{+}}{\min}\,\,N(\widehat{\boldsymbol{\theta}}-\boldsymbol{c})^{\top}\,\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}(\widehat{\boldsymbol{\theta}}-\boldsymbol{c}).$
Under the least favorable null $H_{0}:\boldsymbol{\theta}=\boldsymbol{0}$, it holds,
(4.17) $W_{N}\overset{d}{\longrightarrow}W\sim\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\bigl(\mathbb{R}^{n}_{+}\bigr)^{\circ})=\chi_{n}^{2}-\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+}).$
Moreover, we have $\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\big(\mathbb{R}^{n}_{+}\bigr)^{\circ})=\chi^{2}_{n}-\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+})$, where the equality holds almost surely.
Under the least favorable null, the Wald-type statistic given by $W_{N}$ is the distance of the centered, scaled estimate $\sqrt{N}\widehat{\boldsymbol{\theta}}$ to the closed, convex cone given by the positive orthant $\mathbb{R}^{n}_{+}$. Tests for the opposite sign (e.g., monotonically decreasing or concavity) are obtained by applying the positivity test to the sign-flipped vector $\boldsymbol{\theta}$.
Remark 4.7.
Unlike the usual form of the Wald test, here we have a one-sided test. The test statistic $W_{N}$ measures the projection error under the Mahalanobis distance $\|\cdot\|_{\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}}$, and its limit law describes how far $\boldsymbol{\theta}$ is from the feasibility region, which is given by the composite null hypothesis. Under $H_{0}$, the asymptotic distribution of $W_{N}$ depends on which inequalities are binding for $\boldsymbol{\theta}$. As the entries get more strictly positive (when $\boldsymbol{\theta}$ moves into the interior of $\mathbb{R}^{n}_{+}$), the test statistic $W_{N}$ gets stochastically smaller: the largest (least favorable case) occurs when all the constraints are binding, that is, all the entries of $\boldsymbol{\theta}$ are zero. In particular, for any $\boldsymbol{\theta}\in\mathbb{R}^{n}_{+}$, $\mathbb{P}_{\boldsymbol{\theta}}(W_{N}\geq c)\leq\mathbb{P}_{\boldsymbol{\theta}=\boldsymbol{0}}(W_{N}\geq c)$. So, we calibrate the critical values (or $p$-values) at the least favorable null $\boldsymbol{\theta}=\boldsymbol{0}$.
The asymptotic distribution of $W_{N}$ stated in Theorem 4.6 has the form as given in Theorem 4.4 and hence, to obtain the $p$-values, we need to calculate $\mathbb{P}\left(\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+})\leq c\right)$. In practice, the tail-probability is estimated via a Monte Carlo replication, see Silvapulle and Sen (2001, Section 3.5), such that the test statistic is solved via a non-negative least squares problem, see Appendix E.
5. Numerical experiments
We assess the finite-sample performance of the test statistic using a limit experiment that matches the asymptotic theory in Section 4. Fix a grid size $n$ (number of test points) and a sample size $N$. Under the least-favorable null, we set $\boldsymbol{\theta}=\boldsymbol{0}$ and generate
$\widehat{\boldsymbol{\theta}}\;=\;\boldsymbol{\theta}+\boldsymbol{\Omega}^{1/2}\boldsymbol{Z}/\sqrt{N},\qquad\boldsymbol{Z}\sim\mathcal{\mathcal{N}}(\boldsymbol{0},\boldsymbol{I}_{n}),$
so that $\sqrt{N}\,\widehat{\boldsymbol{\theta}}\overset{d}{\longrightarrow}\mathcal{\mathcal{N}}(\boldsymbol{0},\boldsymbol{\Omega})$ for some positive definite covariance matrix $\boldsymbol{\Omega}$. We compute
$W_{N}=N\,\min_{\boldsymbol{c}\in\mathbb{R}^{n}_{+}}\,\,(\widehat{\boldsymbol{\theta}}-\boldsymbol{c})^{\top}\widehat{\boldsymbol{\Omega}}^{-1}(\widehat{\boldsymbol{\theta}}-\boldsymbol{c}),$
where $\widehat{\boldsymbol{\Omega}}$ is a plug-in covariance that asymptotically converges to $\boldsymbol{\Omega}$. The critical values and $p$-values are obtained via Monte Carlo replications, see Silvapulle and Sen (2001, Section 3.5) for details.
In our experiments, we consider three designs of $\boldsymbol{\Omega}$ (base truth): $(i)$ Identity $\boldsymbol{\Omega}=\boldsymbol{I}_{n}$; $(ii)$ Decaying spectrum $\boldsymbol{\Omega}=\boldsymbol{U}\operatorname{diag}(\boldsymbol{\lambda})\boldsymbol{U}^{\top}$ with decreasing $\lambda_{j}$; $(iii)$ Spiked spectrum (baseline and and spiked eigenvalues with some bulk evenly spaced within a range).
For checking the power robustness of the test, we study three violations whose total signal (in the $\ell_{2}$-norm) is comparable across sparsity levels. Let $k_{\mathrm{mild}}\mathrel{\mathrel{\mathop{:}}=}0.05n$, $k_{\mathrm{mod}}\mathrel{\mathrel{\mathop{:}}=}0.10n$, $k_{\mathrm{strong}}\mathrel{\mathrel{\mathop{:}}=}0.25n$. For “mild” and “moderate” violations we target total signal levels $S_{\mathrm{mild}}=c_{\mathrm{mild}}\sqrt{\log n}$ and $S_{\mathrm{mod}}=c_{\mathrm{mod}}\sqrt{\log n}$, so the per-coordinate shifts are $\delta_{\mathrm{mild}}=S_{\mathrm{mild}}/\sqrt{k_{\mathrm{mild}}}$ and $\delta_{\mathrm{mod}}=S_{\mathrm{mod}}/\sqrt{k_{\mathrm{mod}}}$. For the “strong/dense” violation we set $\delta_{\mathrm{strong}}=S_{\mathrm{strong}}/\sqrt{k_{\mathrm{strong}}}$ with $S_{\mathrm{strong}}=c_{\mathrm{strong}}\sqrt{\log n}$. In each Monte Carlo replication, we select a random support of size $k$ and shift those coordinates by $\delta$ to violate the positivity constraint of $\boldsymbol{\theta}$.
Figure 1. Performance of test statistic: size and power vs. NN. Columns: covariance designs ($\boldsymbol{\Omega}=\boldsymbol{I}_{n}$, SVD-decay, SVD-spike); Rows: grid sizes $n\in\{10,50,100\}$; Curves show empirical size ($H_{0}$) and power under mild/moderate/strong violations with equal-$\ell_{2}$ scaling; dashed line marks the nominal size $\alpha=0.05$.
We vary $(n,N)\in\{10,50,100\}\times\{500,1000,1500,2000\}$. For each scenario, we report the empirical size at $\alpha=0.05$ and power against the three alternatives (mild/moderate/strong). We use between $100$ and $500$ replications per point. The results are reported in Figure 1. We observe that across all covariance designs (identity, SVD-decay, SVD-spike), the procedure exhibits excellent size control: the $H_{0}$ rejection rates (blue) remain close to the nominal $5\%$ line for every $(n,N)$. Power increases monotonically in $N$ and with violation strength, approaching one rapidly for the moderate and strong alternatives, while the mild alternative shows steady gains as $N$ grows. Under equal-$\ell_{2}$ scaling, the behavior is comparable across $n\in\{10,50,100\}$, indicating robustness to grid size and to the spectrum of $\boldsymbol{\Omega}$.
6. Conclusion and future work
We have formulated a nonparametric framework based on RKHS for the mean-variance optimization task, wherein the task functional is linear in both the values of the function and its gradients up to a fixed order. We establish a representer theorem that implies the existence of a finite-dimensional optimal solution to the given empirical problem. Consistency and a functional central limit theorem for the empirical optimizer have been demonstrated, and we have derived finite-sample deviation bounds that shows the impact of regularization. Building upon these results, we have introduced a joint Wald-type test statistic designed to assess shape constraints via positivity of derivative evaluations on a finite grid. Numerical experiments indicate that the test maintains appropriate size control and exhibits increasing power with sample size across diverse covariance structures and varying sparsity patterns.
The findings of this study may be leveraged for examining monotonicity or convexity along a specific covariate direction. Furthermore, the problem formulation and the prosposed methdology appears to be well suited for potential applications in portfolio optimization problems, asset pricing and risk analysis.
Appendix A Proofs for Section 3
Proof of Proposition 3.1.
We start by stating the following inequality without proof:
(A.1) $(a+b)^{t}\leq 2^{t-1}(a^{t}+b^{t})\quad\text{for}\quad t\geq 1.$
Now, $\|\widetilde{\psi}_{1}\|_{\mathcal{H}}=\|\psi_{1}-\mu\|_{\mathcal{H}}=\|\psi_{1}-\mathbb{E}[\psi_{1}]\|_{\mathcal{H}}\leq\|\psi_{1}\|_{\mathcal{H}}+\|\mathbb{E}[\psi_{1}]\|_{\mathcal{H}}$. For any $t\geq 1$,
$\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{t}=\Bigl(\|\psi_{1}\|_{\mathcal{H}}+\|\mathbb{E}[\psi_{1}]\|_{\mathcal{H}}\Bigr)^{t}\leq 2^{t-1}\Bigl(\|\psi_{1}\|_{\mathcal{H}}^{t}+\|\mathbb{E}[\psi_{1}]\|_{\mathcal{H}}^{t}\Bigr)\leq 2^{t-1}\Bigl(\|\psi_{1}\|_{\mathcal{H}}^{t}+\mathbb{E}\|\psi_{1}\|_{\mathcal{H}}^{t}\Bigr),$
where we use (A.1) in the first inequality above. Taking expectations of both sides,
(A.2) $\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{t}\leq 2^{t}\,\mathbb{E}\|\psi_{1}\|^{t}_{\mathcal{H}}.$
Under the assumption $\mathbb{E}\|\psi_{1}\|^{4}_{\mathcal{H}}<\infty$, we have, using (A.2) and Jensen’s inequality,
$\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}\leq 4\mathbb{E}\|\psi_{1}\|_{\mathcal{H}}^{2}\leq 4\left(\mathbb{E}\|\psi_{1}\|_{\mathcal{H}}^{4}\right)^{1/2}<\infty.$
Now, using $\mathcal{C}_{1}=\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}-\mathbb{E}[\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}]$, we have,
$\|\widetilde{\mathcal{C}}_{1}\|_{\operatorname{HS}}\leq\|\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}\|_{\operatorname{HS}}+\|\mathbb{E}[\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}]\|_{\operatorname{HS}}=\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}+\|\mathbb{E}[\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}]\|_{\operatorname{HS}}.$
Squaring both sides and using (A.1) gives us
$\|\widetilde{\mathcal{C}}_{1}\|^{2}_{\operatorname{HS}}\leq 2\Bigl(\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{4}+\|\mathbb{E}[\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}]\|_{\operatorname{HS}}^{2}\Bigr)\leq 2\Bigl(\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{4}+\mathbb{E}\|\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}\|_{\operatorname{HS}}^{2}\Bigr)=2\Bigl(\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{4}+\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{4}\Bigr).$
Taking expectations of both sides and using (A.2),
(A.3) $\mathbb{E}\|\widetilde{\mathcal{C}}_{1}\|_{\operatorname{HS}}^{2}\leq 4\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{4}\leq 64\mathbb{E}\|\psi_{1}\|_{\mathcal{H}}^{4}<\infty.$
Finally, from (3.2),
$\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{\mathcal{C}}_{1})\|_{\mathbb{H}}^{2}=\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}+\mathbb{E}\|\widetilde{\mathcal{C}}_{1}\|_{\operatorname{HS}}^{2}<\infty.$
To prove the second part, we start by noting that any rank-one operator $u\otimes u$ is self-adjoint and positive. $\mathnormal{\Sigma}=\mathbb{E}[\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}]=\mathbb{E}[\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}]$, being the Bochner integral of such form of operators, is self-adjoint and positive as well. Now, we show that $\mathnormal{\Sigma}$ is trace-class by using the cyclical property of the trace, which is a linear map.
$\text{tr}(\mathnormal{\Sigma})=\text{tr}(\mathbb{E}[\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1}])=\mathbb{E}[\text{tr}(\widetilde{\psi}_{1}\otimes\widetilde{\psi}_{1})]=\mathbb{E}[\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}]<\infty.$
This shows that $\mathnormal{\Sigma}\in\mathscr{T}(\mathcal{H})\subset\mathscr{H}\mathscr{S}(\mathcal{H})$, i.e., it is Hilbert-Schmidt. Again, $\widehat{\mathnormal{\Sigma}},\,\widetilde{\mathnormal{\Sigma}}$ are the finite-sum averages of positive, self-adjoint, and Hilbert-Schmidt operators, and hence are as such as well. Finally, $\widetilde{}\mathcal{C}_{i}=\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}-\mathnormal{\Sigma}$ is the difference of self-adjoint and Hilbert-Schmidt operators, and hence follows these properties as well. It also satisfies $\mathbb{E}[\widetilde{}\mathcal{C}_{i}]=\mathbb{E}[\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}]-\mathnormal{\Sigma}=0$, which concludes the proof. ∎
Proof of Proposition 3.2.
We start with
$J_{\lambda}(h)=-\langle h,\mu\rangle_{\mathcal{H}}+\frac{1}{2}\langle h,\mathnormal{\Sigma}h\rangle_{\mathcal{H}}+\frac{\lambda}{2}\|h\|_{\mathcal{H}}^{2}=-\langle h,\mu\rangle_{\mathcal{H}}+\frac{1}{2}\langle h,\mathnormal{\Sigma}_{\lambda}h\rangle_{\mathcal{H}}.$
For any increment $\theta\in\mathcal{H}$,
$J_{\lambda}(h+\theta)-J_{\lambda}(h)$ $=\frac{1}{2}\Bigl(\langle(h+\theta),\mathnormal{\Sigma}_{\lambda}(h+\theta)\rangle_{\mathcal{H}}-\langle h,\mathnormal{\Sigma}_{\lambda}h\rangle_{\mathcal{H}}\Bigr)-\langle\theta,\mu\rangle_{\mathcal{H}}$
$=\langle\theta,\mathnormal{\Sigma}_{\lambda}h-\mu\rangle_{\mathcal{H}}+\frac{1}{2}\langle\theta,\mathnormal{\Sigma}_{\lambda}\theta\rangle_{\mathcal{H}}.$
By the property of the operator norm on $\mathscr{B}(\mathcal{H})$, we have $\|\mathnormal{\Sigma}_{\lambda}\|_{\text{op}}=\|\mathnormal{\Sigma}+\lambda I\|_{\text{op}}\leq\|\mathnormal{\Sigma}\|_{\text{op}}+\lambda<\infty$, since $\mathnormal{\Sigma}\in\mathscr{H}\mathscr{S}(\mathcal{H})\subset\mathscr{B}(\mathcal{H})$. Hence, we have,
$\frac{|J_{\lambda}(h+\theta)-J_{\lambda}(h)-\langle\theta,\mathnormal{\Sigma}_{\lambda}h-\mu\rangle_{\mathcal{H}}|}{\|\theta\|_{\mathcal{H}}}=\frac{|\langle\theta,\mathnormal{\Sigma}_{\lambda}\theta\rangle_{\mathcal{H}}|}{2\|\theta\|_{\mathcal{H}}}\leq\frac{1}{2}\,\|\mathnormal{\Sigma}_{\lambda}\|_{\text{op}}\,\|\theta\|_{\mathcal{H}}\to 0,\quad\text{as}\quad\|\theta\|_{\mathcal{H}}\to 0.$
Hence, $J_{\lambda}$ is Frechét differentiable, see Garling [2014, Chapter 17], with the unique derivative $\nabla J_{\lambda}(h)=\mathnormal{\Sigma}_{\lambda}-\mu$. Therefore, the first-order condition $\nabla J_{\lambda}(h)=0$ implies the normal equation
$\mathnormal{\Sigma}_{\lambda}h_{\lambda}=\mu\iff h_{\lambda}=\mathnormal{\Sigma}_{\lambda}^{-1}\mu,$
since $\mathnormal{\Sigma}_{\lambda}^{-1}$ is well-defined and belongs to $\mathscr{B}(\mathcal{H})$. A similar argument also gives
$\widehat{\mathnormal{\Sigma}}_{\lambda}\widehat{h}_{\lambda}=\widehat{\mu}\iff\widehat{h}_{\lambda}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\widehat{\mu}$,
as $\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\in\mathscr{B}(\mathcal{H})$ is well-defined. This proves the required proposition. ∎
Lemma A.1 (Consistency results).
Let $\mu,\,\mathnormal{\Sigma},\,\widehat{\mu},\,\widehat{\mathnormal{\Sigma}}$ be defined as in (2.9). Then, it holds,
$\|\widehat{\mu}-\mu\|_{\mathcal{H}}\overset{a.s.}{\longrightarrow}0,\qquad\|\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}0.$
Proof.
From (3.5), $(\widehat{\mu}-\mu)$ can be written as the empirical average of the zero-mean i.i.d. vectors $\widetilde{\psi}_{i}$ in the separable Hilbert space $\mathcal{H}$ satisfying $\mathbb{E}\|\widetilde{\psi}_{i}\|_{\mathcal{H}}=\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}<\infty$, since we have assumed the existence of the fourth moment, see Proposition 3.1. The first claim now follows from the strong law of large numbers (SLNN), see Bosq [2000, Theorem 2.4].
For the second claim, we begin by noting from (3.5) that $\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma}$ is the empirical average of zero-mean i.i.d. vectors $\widetilde{\mathcal{C}}_{i}$ in the separable Hilbert space $\mathscr{H}\mathscr{S}(\mathcal{H})$ [^1] For any separable Hilbert space $\mathcal{H}$, the space of Hilbert-Schmidt operators $\mathscr{H}\mathscr{S}(\mathcal{H})$ is also separable. that satisfies $\mathbb{E}\|\widetilde{}\mathcal{C}_{i}\|_{\operatorname{HS}}=\mathbb{E}\|\widetilde{}\mathcal{C}_{1}\|_{\operatorname{HS}}<\infty$; this follows from (A.3). Using the same SLNN as above, we can conclude that $\|\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}0$. Now, we can write $(\psi_{i}-\widehat{\mu})=(\psi_{i}-\mu)-(\widehat{\mu}-\mu)=\widetilde{\psi}_{i}-(\widehat{\mu}-\mu)$. Therefore,
(A.4) $(\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}=-\widetilde{\psi}_{i}\otimes(\widehat{\mu}-\mu)-(\widehat{\mu}-\mu)\otimes\widetilde{\psi}_{i}+(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu).$
Taking the empirical expectation of both sides,
(A.5) $\begin{split}\widehat{\mathnormal{\Sigma}}-\widetilde{\mathnormal{\Sigma}}&=\widehat{}\mathbb{E}[(\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})]-\widehat{}\mathbb{E}[\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i}]\\ &=-\widehat{}\mathbb{E}[\widetilde{\psi}_{i}]\otimes(\widehat{\mu}-\mu)-(\widehat{\mu}-\mu)\otimes\widehat{}\mathbb{E}[\widetilde{\psi}_{i}]+(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)\\ &\overset{(\star)}{=}-(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)-(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)+(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)\\ &=-(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu),\end{split}$
where $(\star)$ follows from (3.5). Hence, $\|\widehat{\mathnormal{\Sigma}}-\widetilde{\mathnormal{\Sigma}}\|_{\operatorname{HS}}=\|-(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)\|_{\operatorname{HS}}=\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}\overset{a.s.}{\longrightarrow}0$ by CMT. So, $\|\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\operatorname{HS}}\leq\|\widehat{\mathnormal{\Sigma}}-\widetilde{\mathnormal{\Sigma}}\|_{\operatorname{HS}}+\|\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}0$, which concludes the proof of the second claim. ∎
Lemma A.2 (Error decomposition).
Consider the expression of $\widehat{h}_{\lambda},\,h_{\lambda}$ as in (3.6). Then,
$\widehat{h}_{\lambda}-h_{\lambda}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}((\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda})+r_{N},\qquad r_{N}\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}(\widetilde{\mathnormal{\Sigma}}-\widehat{\mathnormal{\Sigma}})h_{\lambda}.$
Proof.
Starting from (3.6), we have the following calculation: $\widehat{h}_{\lambda}-h_{\lambda}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\widehat{\mu}-\mathnormal{\Sigma}_{\lambda}^{-1}\mu=\left(\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\widehat{\mu}-\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\mu\right)+\left(\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\mu-\mathnormal{\Sigma}_{\lambda}^{-1}\mu\right)=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}(\widehat{\mu}-\mu)+\left(\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}-\mathnormal{\Sigma}_{\lambda}^{-1}\right)\mu$. Now, for any invertible operators $A,B$, it holds, $A^{-1}-B^{-1}=A^{-1}(B-A)B^{-1}$. Hence,
$\widehat{h}_{\lambda}-h_{\lambda}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}(\widehat{\mu}-\mu)+\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\left(\mathnormal{\Sigma}-\widehat{\mathnormal{\Sigma}}\right)\mathnormal{\Sigma}_{\lambda}^{-1}\mu=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}(\widehat{\mu}-\mu)+\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\left(\mathnormal{\Sigma}-\widehat{\mathnormal{\Sigma}}\right)h_{\lambda}.$
Defining $r_{N}\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}(\widetilde{\mathnormal{\Sigma}}-\widehat{\mathnormal{\Sigma}})h_{\lambda}$, we obtain,
$\widehat{h}_{\lambda}-h_{\lambda}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\left((\widehat{\mu}-\mu)-(\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda}\right)=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\left((\widehat{\mu}-\mu)-(\widetilde{\mathnormal{\Sigma}}-\mathnormal{\Sigma})h_{\lambda}\right)+r_{N},$
which proves the lemma. ∎
Lemma A.3 (Properties of $\widehat{\mu}$).
Let $\widehat{\mu},\,\mu$ be defined as in (2.9). If $\mathbb{E}\|{\psi}_{1}\|_{\mathcal{H}}^{2}<\infty$, then
$\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}=o_{\mathbb{P}}(N^{-1/2}).$
Moreover, for any $\delta\in(0,1)$,
$\mathbb{P}\left(\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}>\frac{2\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}}{N\delta}\right)\leq\frac{\delta}{2}.$
Proof.
Since $\widetilde{\psi}_{i}$ are zero-mean i.i.d. $\mathcal{H}$-valued random vectors with finite second moments,
$\mathbb{E}\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}=\frac{1}{N^{2}}\mathbb{E}\|\sum_{i=1}^{N}\widetilde{\psi}_{i}\|_{\mathcal{H}}^{2}=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathbb{E}\|\widetilde{\psi}_{i}\|_{\mathcal{H}}^{2}=\frac{1}{N}\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}.$
The second equality is due to the implication of weak orthogonality from the independence of zero-mean random vectors, see Bosq [2000, Definition 1.2]. By Markov’s inequality applied to $\sqrt{N}\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}$, for any $\varepsilon>0$,
$\mathbb{P}\left(\sqrt{N}\,\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}>\varepsilon\right)\leq\frac{\sqrt{N}\,\mathbb{E}\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}}{\varepsilon}=\frac{1}{\sqrt{N}}\cdot\frac{\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}}{\varepsilon}.$
Therefore, $\underset{N\to\infty}{\lim}\mathbb{P}(\sqrt{N}\,\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}>\varepsilon)=0$, which implies $\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}=o_{\mathbb{P}}(N^{-1/2})$.
For the second part of the claim, for any $\delta\in(0,1)$, we apply Markov’s inequality directly to $\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}\geq 0$ to obtain
$\mathbb{P}\left(\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}>\frac{2\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}}{N\delta}\right)\leq\mathbb{E}\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}\cdot\frac{N\delta}{2\mathbb{E}\|\widetilde{\psi}_{1}\|_{\mathcal{H}}^{2}}=\frac{\delta}{2}.$
Lemma A.4 (Remainder term).
Let $r_{N}$ be defined as in (3.7). Then, $\sqrt{N}\,r_{N}\overset{\mathbb{P}}{\longrightarrow}0$.
Proof.
From the definition of $r_{N}$ in (3.7) and the expression in (A.5), we have
$\|r_{N}\|_{\mathcal{H}}=\|\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}(\widetilde{\mathnormal{\Sigma}}-\widehat{\mathnormal{\Sigma}})h_{\lambda}\|_{\mathcal{H}}\leq\|\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\|_{\text{op}}\,\|(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)\,h_{\lambda}\|_{\mathcal{H}}.$
Since $\widehat{\mathnormal{\Sigma}}$ is a positive operator and $\lambda>0,\|\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\|_{\text{op}}=\lambda_{\text{max}}(\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1})=1/(\lambda_{\text{min}}(\widehat{\mathnormal{\Sigma}}_{\lambda}))\leq 1/\lambda$. Again,
$\|(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)\,h_{\lambda}\|_{\mathcal{H}}$ $\leq\|(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)\|_{\text{op}}\,\|h_{\lambda}\|_{\mathcal{H}}$
$\leq\|(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)\|_{\operatorname{HS}}\,\|h_{\lambda}\|_{\mathcal{H}}=\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}\,\|h_{\lambda}\|_{\mathcal{H}}.$
Therefore,
(A.6) $\|r_{N}\|_{\mathcal{H}}\leq\frac{\|h_{\lambda}\|_{\mathcal{H}}}{\lambda}\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}.$
As $0<\|h_{\lambda}\|_{\mathcal{H}}/\lambda<\infty$ and $\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}=o_{\mathbb{P}}(N^{-1/2})$ from Lemma A.3, the conclusion follows. ∎
Lemma A.5 (Properties of $F$).
Let $F\colon\mathbb{H}\to\mathcal{H}$ be defined as $F(h,\mathcal{C})\mathrel{\mathrel{\mathop{:}}=}h-\mathcal{C}h_{\lambda}$. Then, $F$ is a linear and bounded map that satisfies
$\|F\|_{\operatorname{op}}\leq\sqrt{1+\|h_{\lambda}\|_{\mathcal{H}}^{2}}<\infty.$
Proof.
$F$ is linear by construction. We now show that it is bounded in the operator norm. Starting from the definition of the operator norm,
$\|F\|_{\text{op}}=\underset{\|(h,\mathcal{C})\|_{\mathbb{H}}=1}{\operatorname{sup}}\|h-\mathcal{C}h_{\lambda}\|_{\mathcal{H}}\leq\underset{\|(h,\mathcal{C})\|_{\mathbb{H}}=1}{\operatorname{sup}}\|h\|_{\mathcal{H}}+\|\mathcal{C}h_{\lambda}\|_{\mathcal{H}}\leq\underset{\|(h,\mathcal{C})\|_{\mathbb{H}}=1}{\operatorname{sup}}\|h\|_{\mathcal{H}}+\|\mathcal{C}\|_{\text{op}}\|h_{\lambda}\|_{\mathcal{H}}.$
Using the inequality $a+bc\leq\sqrt{a^{2}+b^{2}}\,\sqrt{1+c^{2}}$ (this is due to Cauchy-Schwarz) applied to the expression above, we obtain
$\|F\|_{\text{op}}$ $\leq\underset{\|(h,\mathcal{C})\|_{\mathbb{H}}=1}{\operatorname{sup}}\left\{\sqrt{\|h\|_{\mathcal{H}}^{2}+\|\mathcal{C}\|_{\operatorname{HS}}^{2}}\cdot\sqrt{1+\|h_{\lambda}\|_{\mathcal{H}}^{2}}\right\}$
$=\underset{\|(h,\mathcal{C})\|_{\mathbb{H}}=1}{\operatorname{sup}}\left\{\|(h,\mathcal{C})\|_{\mathbb{H}}\cdot\sqrt{1+\|h_{\lambda}\|_{\mathcal{H}}^{2}}\right\}=\sqrt{1+\|h_{\lambda}\|_{\mathcal{H}}^{2}}<\infty.$
Lemma A.6 (Finite-sample bound for $\bar{S}_{N}$).
Let $\bar{S}_{N}=\frac{1}{N}\sum_{i=1}^{N}(\widetilde{\psi}_{i},\widetilde{}\mathcal{C}_{i})$ where $\widetilde{\psi}_{i},\,\widetilde{}\mathcal{C}_{i}$ as defined in Section 3. Under the assumptions of Proposition 3.1, it holds for any $\delta\in(0,1)$,
(A.7) $\mathbb{P}\left(\|\bar{S}_{N}\|_{\mathbb{H}}>\sqrt{\frac{2\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}}{N\delta}}\right)\leq\frac{\delta}{2}.$
Proof.
Since $(\widetilde{\psi}_{i},\widetilde{}\mathcal{C}_{i})$ are zero-mean i.i.d. random vectors in the separable Hilbert space $\mathbb{H}$,
$\mathbb{E}\|\bar{S}_{N}\|_{\mathbb{H}}^{2}=\frac{1}{N^{2}}\mathbb{E}\|\sum_{i=1}^{N}(\widetilde{\psi}_{i},\widetilde{}\mathcal{C}_{i})\|_{\mathbb{H}}^{2}=\frac{1}{N^{2}}\sum_{i=1}^{N}\mathbb{E}\|(\widetilde{\psi}_{i},\widetilde{}\mathcal{C}_{i})\|_{\mathbb{H}}^{2}=\frac{1}{N}\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}.$
The second equality follows since zero-mean i.i.d. vectors in a separable Hilbert space imply weak orthogonality, see Bosq [2000, Definition 1.2]. Now, applying Markov’s inequality, for any $\varepsilon>0$,
$\mathbb{P}\left(\|\bar{S}_{N}\|_{\mathbb{H}}>\varepsilon\right)=\mathbb{P}\left(\|\bar{S}_{N}\|_{\mathbb{H}}^{2}>\varepsilon^{2}\right)\leq\frac{\mathbb{E}\|\bar{S}_{N}\|_{\mathbb{H}}^{2}}{\varepsilon^{2}}=\frac{\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}}{N\varepsilon^{2}}.$
For any $\delta\in(0,1)$, choosing $\varepsilon=\sqrt{\frac{2\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}}{N\delta}}$ gives the required inequality. ∎
Appendix B Lemmas for Section 4
Lemma B.1 (Asymptotic convergence of mean-squared error).
Define $\Delta_{i}\mathrel{\mathrel{\mathop{:}}=}\widehat{F}_{i}-F_{i}$. Under the conditions of Proposition 3.1, it holds,
(B.1) $\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}^{2}\overset{a.s.}{\longrightarrow}0.$
Proof.
From the definition of $\Delta_{i}=\widehat{F}_{i}-F_{i}$, we obtain
$\Delta_{i}$ $=(\psi_{i}-\widehat{\mu})-\left((\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-\widehat{\mathnormal{\Sigma}}\right)\widehat{h}_{\lambda}-\left[(\psi_{i}-\mu)-\Bigl((\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\Bigr)h_{\lambda}\right]$
$=(\mu-\widehat{\mu})-\Bigl((\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)\Bigr)\widehat{h}_{\lambda}$
$\qquad-\Bigl((\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\Bigr)(\widehat{h}_{\lambda}-h_{\lambda})+(\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma})\widehat{h}_{\lambda}.$
Therefore, we obtain the following:
$\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}^{2}$ $\leq\underbrace{\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}}_{(I)}+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\|\bigl((\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)\bigr)\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}}_{(II)}$
$\qquad+\underbrace{\frac{1}{N}\sum_{i=1}^{N}\|\bigl((\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\bigr)(\widehat{h}_{\lambda}-h_{\lambda})\|_{\mathcal{H}}^{2}}_{(III)}+\underbrace{\|(\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma})\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}}_{(IV)}.$
We now show the convergence for each of these quantities. From $(i)$ of(A.1), we have
(B.2) $(I)=\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2}\overset{a.s.}{\longrightarrow}0.$
To show that $(II)\overset{a.s.}{\longrightarrow}0$, first note that
$(II)$ $=\frac{1}{N}\sum_{i=1}^{N}\|\bigl((\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)\bigr)\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}$
$\leq\frac{\|\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}}{N}\sum_{i=1}^{N}\|(\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)\|_{\text{op}}^{2}$
$\leq\frac{\|\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}}{N}\sum_{i=1}^{N}\|(\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)\|_{\operatorname{HS}}^{2}.$
Using the decomposition in (A.4) with $\widetilde{\psi}_{i}=\psi_{i}-\mu$, we obtain that
$(II)$ $\leq\|\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}\,\Bigg(\frac{1}{N}\sum_{i=1}^{N}\|\widetilde{\psi}_{i}\otimes(\widehat{\mu}-\mu)\|^{2}_{\operatorname{HS}}+\frac{1}{N}\sum_{i=1}^{N}\|(\widehat{\mu}-\mu)\otimes\widetilde{\psi}_{i}\|_{\operatorname{HS}}^{2}$
$\qquad\qquad+\frac{1}{N}\sum_{i=1}^{N}\|(\widehat{\mu}-\mu)\otimes(\widehat{\mu}-\mu)\|_{\operatorname{HS}}^{2}\Bigg)$
$=\|\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}\left(\frac{2\|\widehat{\mu}-\mu\|^{2}_{\mathcal{H}}}{N}\sum_{i=1}^{N}\|\widetilde{\psi}_{i}\|_{\mathcal{H}}^{2}+\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{4}\right),$
where we used $\|f\otimes g\|_{\operatorname{HS}}=\|f\|_{\mathcal{H}}\,\|g\|_{\mathcal{H}}$ in the last equality. Now, under the conditions of Proposition 3.1, $\frac{1}{N}\sum_{i=1}^{N}\|\widetilde{\psi}_{i}\|_{\mathcal{H}}^{2}\overset{a.s.}{\longrightarrow}\mathbb{E}\|\widetilde{\psi}_{i}\|^{2}_{\mathcal{H}}=\mathbb{E}\|\widetilde{\psi}_{1}\|^{2}_{\mathcal{H}}<\infty$, from the SLNN. Applying CMT to the result in $(i)$ of Lemma A.1, we have, $\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{2},\,\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{4}\overset{a.s.}{\longrightarrow}0$. Combining these and using that $\|\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}<\infty$, we can conclude:
(B.3) $(II)\leq\|\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}\left(\frac{2\|\widehat{\mu}-\mu\|^{2}_{\mathcal{H}}}{N}\sum_{i=1}^{N}\|\widetilde{\psi}_{i}\|_{\mathcal{H}}^{2}+\|\widehat{\mu}-\mu\|_{\mathcal{H}}^{4}\right)\overset{a.s.}{\longrightarrow}0.$
For the third term, we start with the decomposition,
$(III)$ $=\frac{1}{N}\sum_{i=1}^{N}\|\bigl((\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\bigr)(\widehat{h}_{\lambda}-h_{\lambda})\|_{\mathcal{H}}^{2}$
$\leq\frac{\|\widehat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}^{2}}{N}\sum_{i=1}^{N}\|(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\|_{\text{op}}$
$\leq\frac{\|\widehat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}^{2}}{N}\sum_{i=1}^{N}\|(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\|_{\operatorname{HS}}.$
From the defintion of $\widetilde{}\mathcal{C}_{i}=(\widetilde{\psi}_{i}\otimes\widetilde{\psi}_{i})-\mathnormal{\Sigma}$ from (3.3), we obtain from the SLN:
$\frac{1}{N}\sum_{i=1}^{N}\|(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\|_{\operatorname{HS}}=\frac{1}{N}\sum_{i=1}^{N}\|\widetilde{\mathcal{C}}_{i}\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}\mathbb{E}\|\widetilde{C}_{i}\|_{\operatorname{HS}}^{2}=\mathbb{E}\|\widetilde{C}_{1}\|_{\operatorname{HS}}^{2}<\infty.$
From $(i)$ of Proposition 3.4, $\|\widehat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}^{2}\overset{a.s.}{\longrightarrow}0$. Hence,
(B.4) $(III)\leq\frac{\|\widehat{h}_{\lambda}-h_{\lambda}\|_{\mathcal{H}}^{2}}{N}\sum_{i=1}^{N}\|(\psi_{i}-\mu)\otimes(\psi_{i}-\mu)-\mathnormal{\Sigma}\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}0.$
Finally, from $(ii)$ of Lemma A.1, $\|\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\operatorname{HS}}^{2}\overset{a.s.}{\longrightarrow}0$. So, the fourth term satisfies
(B.5) $(IV)=\|(\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma})\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}\leq\|\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\text{op}}^{2}\,\|\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}\leq\|\widehat{\mathnormal{\Sigma}}-\mathnormal{\Sigma}\|_{\operatorname{HS}}^{2}\,\|\widehat{h}_{\lambda}\|_{\mathcal{H}}^{2}\overset{a.s.}{\longrightarrow}0.$
Combining the results of convergence for the individual terms, the conclusion follows. ∎
Lemma B.2 ((Asymptotic) boundedness).
Under the conditions of Proposition 3.1, it holds,
(B.6) $\|u_{j}\|_{\mathcal{H}}<\infty,\qquad\|\widehat{u}_{j}\|_{\mathcal{H}}<\infty\quad\text{for all}\quad 1\leq j\leq n.$
Moreover,
(B.7) $\|\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}<\infty,\qquad\|\widehat{}\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}=\mathcal{O}(1)\;\text{almost surely as}\;N\to\infty.$
Proof.
For any $j=1,\ldots,n$, we have the calculation:
$\|u_{j}\|_{\mathcal{H}}=\|\mathnormal{\Sigma}_{\lambda}^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\|_{\mathcal{H}}\leq\|\mathnormal{\Sigma}_{\lambda}^{-1}\|_{\text{op}}\,\|\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\|_{\mathcal{H}}\leq\frac{\|\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\|_{\mathcal{H}}}{\lambda}<\infty,$
since $\|\mathnormal{\Sigma}_{\lambda}\|_{\text{op}}=\lambda_{\max}(\mathnormal{\Sigma}_{\lambda}^{-1})=1/(\lambda_{\text{min}}(\mathnormal{\Sigma}_{\lambda}))\leq 1/\lambda$ as $\mathnormal{\Sigma}$ is a positive operator. A similar calculation ensures that $\|\widehat{u}_{j}\|_{\mathcal{H}}\leq\frac{\|\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\|_{\mathcal{H}}}{\lambda}<\infty$ for all $j=1,\ldots,n$.
To show $\|\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}<\infty$, we start from the definition of $F_{i}$, see (4.4). Then,
(B.8) $\mathbb{E}\|F_{1}\|_{\mathcal{H}}^{2}=\mathbb{E}\|F(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathcal{H}}^{2}\leq\|F\|_{\text{op}}^{2}\;\mathbb{E}\|(\widetilde{\psi}_{1},\widetilde{}\mathcal{C}_{1})\|_{\mathbb{H}}^{2}<\infty,$
which follows from Proposition 3.1 and Lemma A.5. Therefore, using the above inequality,
(B.9) $\|\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}=\|\mathbb{E}[F_{i}\otimes F_{i}]\|_{\operatorname{HS}}=\|\mathbb{E}[F_{1}\otimes F_{1}]\|_{\operatorname{HS}}\leq\mathbb{E}\|F_{1}\otimes F_{1}\|_{\operatorname{HS}}=\mathbb{E}\|F_{1}\|^{2}_{\mathcal{H}}<\infty.$
To show the final result, we start by writing $\widehat{F}_{i}=\Delta_{i}+F_{i}$, where $\Delta_{i}\mathrel{\mathrel{\mathop{:}}=}\widehat{F}_{i}-F_{i}$, such that
$\frac{1}{N}\sum_{i=1}^{N}\|\widehat{F}_{i}\|_{\mathcal{H}}^{2}\leq\underbrace{\frac{2}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}^{2}}_{\overset{a.s.}{\longrightarrow}0}+\underbrace{\frac{2}{N}\sum_{i=1}^{N}\|F_{i}\|_{\mathcal{H}}^{2}}_{\overset{a.s.}{\longrightarrow}2\mathbb{E}\|F_{1}\|_{\mathcal{H}}^{2}}\overset{a.s.}{\longrightarrow}2\mathbb{E}\|F_{1}\|_{\mathcal{H}}^{2}<\infty,$
where we use Lemma B.1, and SLNN since (B.8) holds true. Therefore, we can conclude that:
(B.10) $\frac{1}{N}\sum_{i=1}^{N}\|\widehat{F}_{i}\|_{\mathcal{H}}^{2}=\mathcal{O}(1)\;\text{almost surely as}\;N\to\infty.$
Using (B.10) along with the definition of $\widehat{}\mathcal{Q}_{\lambda}$ from (4.3), we get for $N\to\infty$,
$\|\widehat{}\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}=\Bigl\|\frac{1}{N}\sum_{i=1}^{N}\widehat{F}_{i}\otimes\widehat{F}_{i}\Bigr\|_{\operatorname{HS}}\leq\frac{1}{N}\sum_{i=1}^{N}\|\widehat{F}_{i}\otimes\widehat{F}_{i}\|_{\operatorname{HS}}=\frac{1}{N}\sum_{i=1}^{N}\|\widehat{F}_{i}\|_{\mathcal{H}}^{2}=\mathcal{O}(1)\;\text{almost surely.}$
Lemma B.3 (Consistency of $\widehat{\mathcal{Q}}_{\lambda}$).
Let $\widehat{\mathcal{Q}}_{\lambda}$ and $\mathcal{Q}_{\lambda}$ be defined as in (4.3) and (3.10) respectively. Under the conditions of Proposition 3.1, it holds,
$\|\widehat{\mathcal{Q}}_{\lambda}-\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}0.$
Proof.
We start with
$\widehat{}\mathcal{Q}_{\lambda}-\mathcal{Q}_{\lambda}$ $=\frac{1}{N}\sum_{i=1}^{N}\widehat{F}_{i}\otimes\widehat{F}_{i}-\mathbb{E}[F_{i}\otimes F_{i}]$
$=\frac{1}{N}\sum_{i=1}^{N}\Bigl(\widehat{F}_{i}\otimes\widehat{F}_{i}-F_{i}\otimes F_{i}\Bigr)+\frac{1}{N}\sum_{i=1}^{N}F_{i}\otimes F_{i}-\mathbb{E}[F_{i}\otimes F_{i}].$
Therefore,
$\|\widehat{}\mathcal{Q}_{\lambda}-\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}\leq\underbrace{\Bigl\|\frac{1}{N}\sum_{i=1}^{N}(\widehat{F}_{i}\otimes\widehat{F}_{i}-F_{i}\otimes F_{i})\Bigr\|_{\operatorname{HS}}}_{(I)}+\underbrace{\Bigl\|\frac{1}{N}\sum_{i=1}^{N}F_{i}\otimes F_{i}-\mathbb{E}[F_{i}\otimes F_{i}]\Bigr\|_{\operatorname{HS}}}_{(II)}.$
With $\Delta_{i}\mathrel{\mathrel{\mathop{:}}=}\widehat{F}_{i}-F_{i}$, we can write $\widehat{F}_{i}\otimes\widehat{F}_{i}-F_{i}\otimes F_{i}=\Delta_{i}\otimes\widehat{F}_{i}+F_{i}\otimes\Delta_{i}$. Hence,
$(I)\leq\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\otimes\widehat{F}_{i}+F_{i}\otimes\Delta_{i}\|_{\operatorname{HS}}$ $\leq\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\otimes\widehat{F}_{i}\|_{\operatorname{HS}}+\frac{1}{N}\sum_{i=1}^{N}\|F_{i}\otimes\Delta_{i}\|_{\operatorname{HS}}$
$=\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}\,\|\widehat{F}_{i}\|_{\mathcal{H}}+\frac{1}{N}\sum_{i=1}^{N}\|F_{i}\|_{\mathcal{H}}\,\|\Delta_{i}\|_{\mathcal{H}}.$
From the Cauchy-Schwarz inequality,
$\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}\,\|\widehat{F}_{i}\|_{\mathcal{H}}$ $\leq\left(\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}^{2}\right)^{1/2}\,\left(\frac{1}{N}\sum_{i=1}^{N}\|\widehat{F}_{i}\|_{\mathcal{H}}^{2}\right)^{1/2}$
$\frac{1}{N}\sum_{i=1}^{N}\|F_{i}\|_{\mathcal{H}}\,\|\Delta_{i}\|_{\mathcal{H}}$ $\leq\left(\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}^{2}\right)^{1/2}\,\left(\frac{1}{N}\sum_{i=1}^{N}\|F_{i}\|_{\mathcal{H}}^{2}\right)^{1/2}.$
From Lemma B.1, $\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}^{2}\overset{a.s.}{\longrightarrow}0$, while $\frac{1}{N}\sum_{i=1}^{N}\|\widehat{F}_{i}\|_{\mathcal{H}}^{2}=\mathcal{O}(1)\;\text{almost surely as}\;N\to\infty$, see (B.10). By SLNN, $\frac{1}{N}\sum_{i=1}^{N}\|F_{i}\|_{\mathcal{H}}^{2}\overset{a.s.}{\longrightarrow}\mathbb{E}\|F_{1}\|_{\mathcal{H}}^{2}<\infty$, see (B.8). Therefore,
$\begin{split}(I)\leq\underbrace{\left(\frac{1}{N}\sum_{i=1}^{N}\|\Delta_{i}\|_{\mathcal{H}}^{2}\right)^{1/2}}_{\overset{a.s.}{\longrightarrow}0}\Bigg(\underbrace{\left(\frac{1}{N}\sum_{i=1}^{N}\|\widehat{F}_{i}\|_{\mathcal{H}}^{2}\right)^{1/2}}_{=\mathcal{O}(1)\;\text{a.s. for}\;N\to\infty}+\underbrace{\left(\frac{1}{N}\sum_{i=1}^{N}\|F_{i}\|_{\mathcal{H}}^{2}\right)^{1/2}}_{\overset{a.s.}{\longrightarrow}\left(\mathbb{E}\|F_{1}\|_{\mathcal{H}}^{2}\right)^{1/2}<\infty}\Bigg)\overset{a.s.}{\longrightarrow}0.\end{split}$
From (B.9) it holds, $\|\mathbb{E}[F_{i}\otimes F_{i}]\|_{\operatorname{HS}}=\mathbb{E}\|F_{1}\|_{\mathcal{H}}^{2}<\infty$. By SLNN,
$(II)=\Bigl\|\frac{1}{N}\sum_{i=1}^{N}F_{i}\otimes F_{i}-\mathbb{E}[F_{i}\otimes F_{i}]\Bigr\|_{\operatorname{HS}}=\Bigl\|\frac{1}{N}\sum_{i=1}^{N}\big(F_{i}\otimes F_{i}-\mathbb{E}[F_{i}\otimes F_{i}]\big)\Bigr\|_{\operatorname{HS}}\overset{a.s.}{\longrightarrow}0.$
Combining the above, we obtain
$\|\widehat{}\mathcal{Q}_{\lambda}-\mathcal{Q}_{\lambda}\|_{\operatorname{HS}}\leq(I)+(II)\overset{a.s.}{\longrightarrow}0.$
which proves the claim. ∎
Lemma B.4 (Action of $\widehat{\mathnormal{\Sigma}}$).
Let $\widetilde{\boldsymbol{\Psi}}$ be defined as in (E.5). Then, $\widehat{\mathnormal{\Sigma}}=\frac{1}{N}\widetilde{\boldsymbol{\Psi}}\widetilde{\boldsymbol{\Psi}}^{\ast}$.
Proof.
[…]
$\widetilde{\boldsymbol{\Psi}}\widetilde{\boldsymbol{\Psi}}^{\ast}(h)=\widetilde{\boldsymbol{\Psi}}\left([\langle h,\psi_{i}-\widehat{\mu}\rangle_{\mathcal{H}}]_{i=1}^{N}\right)=\sum_{i=1}^{N}\langle h,\psi_{i}-\widehat{\mu}\rangle_{\mathcal{H}}(\psi_{i}-\widehat{\mu})=\sum_{i=1}^{N}\left((\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})\right)h.$
Dividing by $N$ then gives the necessary conclusion. ∎
Lemma B.5 (Computation of $\widehat{u}_{j}$).
Let $\widehat{u}_{j}$ be defined as in (4.2). Then, it holds,
$\widehat{u}_{j}=\frac{1}{\lambda}(\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})-\boldsymbol{\Psi}\boldsymbol{\gamma}_{j}),\qquad\boldsymbol{\gamma}_{j}=\frac{1}{N}\boldsymbol{H}\left(\lambda\boldsymbol{I}_{N}+\frac{1}{N}\boldsymbol{H}\boldsymbol{G}\boldsymbol{H}\right)^{-1}[\widetilde{\boldsymbol{G}}_{\mathcal{G}}]_{j},$
where $\boldsymbol{\Psi},\,\boldsymbol{H},\,\boldsymbol{G},\,\boldsymbol{G}_{\mathcal{G}}$ are defined in Section E.
Proof.
We start with $\frac{1}{N}\widetilde{\boldsymbol{\Psi}}^{\ast}\widetilde{\boldsymbol{\Psi}}=\frac{1}{N}\boldsymbol{H}\boldsymbol{\Psi}^{\ast}\boldsymbol{\Psi}\boldsymbol{H}=\frac{1}{N}\boldsymbol{H}\boldsymbol{G}\boldsymbol{H}$. This holds since we can interpret $\boldsymbol{\Psi}$ as the map from $\mathbb{R}^{N}$ to $\mathcal{H}$ acting as $\boldsymbol{\alpha}\mapsto\sum_{i=1}^{N}\alpha_{i}\psi_{i}$ whose adjoint acts as $\boldsymbol{\Psi}^{\ast}(\cdot)=[\langle\cdot,\psi_{i}\rangle_{\mathcal{H}}]_{i=1}^{N}$; hence, $\boldsymbol{\Psi}^{\ast}\boldsymbol{\Psi}=\boldsymbol{\Psi}^{\ast}([\psi_{1},\ldots,\psi_{N}])=\langle\boldsymbol{\Psi}^{\top},\boldsymbol{\Psi}\rangle_{\mathcal{H}}=\boldsymbol{G}$. We now employ the Woodbury identity:
$\left(\frac{1}{N}\widetilde{\boldsymbol{\Psi}}\widetilde{\boldsymbol{\Psi}}^{\ast}+\lambda I\right)^{-1}$ $=\frac{1}{\lambda}\left(\frac{1}{\lambda N}\widetilde{\boldsymbol{\Psi}}\widetilde{\boldsymbol{\Psi}}^{\ast}+I\right)^{-1}$
[…]
Using (E.9) and Lemma B.4, we can write:
$\widehat{u}_{j}=\left(\frac{1}{N}\widetilde{\boldsymbol{\Psi}}\widetilde{\boldsymbol{\Psi}}^{\ast}+\lambda I\right)^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})=\frac{1}{\lambda}\left(I-\frac{1}{N}\widetilde{\boldsymbol{\Psi}}\Bigl(\lambda\boldsymbol{I}_{N}+\frac{1}{N}\boldsymbol{H}\boldsymbol{G}\boldsymbol{H}\Bigr)^{-1}\widetilde{\boldsymbol{\Psi}}^{\ast}\right)\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j}),$
which implies
$\widehat{u}_{j}=\frac{1}{\lambda}\Big(\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})-\frac{1}{N}\boldsymbol{\Psi}\boldsymbol{H}\Bigl(\lambda\boldsymbol{I}_{N}+\frac{1}{N}\boldsymbol{H}\boldsymbol{G}\boldsymbol{H}\Bigr)^{-1}\boldsymbol{H}\boldsymbol{\Psi}^{\ast}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\Big)=\frac{1}{\lambda}\left(\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})-\boldsymbol{\Psi}\boldsymbol{\gamma}_{j}\right),$
where $\boldsymbol{\gamma}_{j}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{N}\boldsymbol{H}\left(\lambda\boldsymbol{I}_{N}+\frac{1}{N}\boldsymbol{H}\boldsymbol{G}\boldsymbol{H}\right)^{-1}\boldsymbol{H}\boldsymbol{\Psi}^{\ast}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})$. Finally, by noting that
$\boldsymbol{H}\boldsymbol{\Psi}^{\ast}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})=[\widetilde{\boldsymbol{\Psi}}^{\ast}\boldsymbol{\Phi}_{\mathcal{G}}]_{j}=[\langle\widetilde{\boldsymbol{\Psi}}^{\top},\boldsymbol{\Phi}_{\mathcal{G}}\rangle_{\mathcal{H}}]_{j}=[\widetilde{\boldsymbol{G}}_{\mathcal{G}}]_{j},$
the claim follows. ∎
Lemma B.6 (Computation of $\boldsymbol{B}$).
Define the matrix $\boldsymbol{B}\mathrel{\mathrel{\mathop{:}}=}\left[\langle\widehat{F}_{i},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}\right]_{i,j=1}^{N,n}\in\mathbb{R}^{N\times n}$, where $\widehat{F}_{i}$ is defined as in (4.3). Then, it holds,
$\boldsymbol{B}=\left(\boldsymbol{I}-\operatorname{diag}(\widetilde{\boldsymbol{h}})\right)\widetilde{\boldsymbol{G}}_{\mathcal{G}}+\frac{1}{N}\boldsymbol{1}\left(\widetilde{\boldsymbol{h}}^{\top}\widetilde{\boldsymbol{G}}_{\mathcal{G}}\right)\in\mathbb{R}^{N\times n}.$
Proof.
Define the pairwise entries of the matrix $\boldsymbol{B}$ as $\beta_{i,j}\mathrel{\mathrel{\mathop{:}}=}\langle\widehat{F}_{i},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}$. From the expression of $\widehat{F}_{i}$ in (4.3), we obtain
$\beta_{i,j}$ $=\left\langle(\psi_{i}-\widehat{\mu})-\left((\psi_{i}-\widehat{\mu})\otimes(\psi_{i}-\widehat{\mu})-\widehat{\mathnormal{\Sigma}}\right)\widehat{h}_{\lambda},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\right\rangle_{\mathcal{H}}$
$=\langle\psi_{i}-\widehat{\mu},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}-\langle\psi_{i}-\widehat{\mu},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}\langle\psi_{i}-\widehat{\mu},\widehat{h}_{\lambda}\rangle_{\mathcal{H}}$
$\qquad+\frac{1}{N}\sum_{i=1}^{N}\langle\psi_{i}-\widehat{\mu},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}\langle\psi_{i}-\widehat{\mu},\widehat{h}_{\lambda}\rangle_{\mathcal{H}}.$
From (E.7), the matrix $\widetilde{\boldsymbol{G}}_{\mathcal{G}}$ contains the parwise entries $\langle\psi_{i}-\widehat{\mu},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}$, while the entries $\langle\psi_{i}-\widehat{\mu},\widehat{h}_{\lambda}\rangle_{\mathcal{H}}$ are encoded in the vector $\widetilde{\boldsymbol{h}}$ from (E.8). Finally, the column means are given the entries of the row vector $\frac{1}{N}\widetilde{\boldsymbol{h}}^{\top}\widetilde{\boldsymbol{G}}_{\mathcal{G}}$. Hence, the matrix form is justified. ∎
Lemma B.7 (Computation of $\widehat{\boldsymbol{\Omega}}_{\lambda}$).
Let $\widehat{\boldsymbol{\Omega}}_{\lambda}$ be defined as in (4.5). Then,
$\widehat{\boldsymbol{\Omega}}_{\lambda}=\frac{1}{N}\boldsymbol{S}^{\top}\boldsymbol{S},\qquad\boldsymbol{S}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{\lambda}(\boldsymbol{B}-\boldsymbol{V}^{\top}\boldsymbol{\Lambda}),$
where
(B.11) $\begin{split}\boldsymbol{B}&\mathrel{\mathrel{\mathop{:}}=}\left[\langle\widehat{F}_{i},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}\right]_{i,j=1}^{N,n}=\left(\boldsymbol{I}-\operatorname{diag}(\widetilde{\boldsymbol{h}})\right)\widetilde{\boldsymbol{G}}_{\mathcal{G}}+\frac{1}{N}\boldsymbol{1}\left(\widetilde{\boldsymbol{h}}^{\top}\widetilde{\boldsymbol{G}}_{\mathcal{G}}\right)\in\mathbb{R}^{N\times n},\\ \boldsymbol{V}&\mathrel{\mathrel{\mathop{:}}=}\left[\boldsymbol{\Psi}^{\ast}\widehat{F}_{1},\ldots,\boldsymbol{\Psi}^{\ast}\widehat{F}_{N}\right]=\left(\boldsymbol{I}-\operatorname{diag}(\widetilde{\boldsymbol{h}})\right)\widetilde{\boldsymbol{G}}+\frac{1}{N}\boldsymbol{1}\left(\widetilde{\boldsymbol{h}}^{\top}\widetilde{\boldsymbol{G}}\right)\in\mathbb{R}^{N\times N},\\ \boldsymbol{\Lambda}&\mathrel{\mathrel{\mathop{:}}=}[\boldsymbol{\gamma}_{1},\ldots,\boldsymbol{\gamma}_{n}]=\frac{1}{N}\boldsymbol{H}\left(\lambda\boldsymbol{I}_{N}+\frac{1}{N}\boldsymbol{H}\boldsymbol{G}\boldsymbol{H}\right)^{-1}\widetilde{\boldsymbol{G}}_{\mathcal{G}}\in\mathbb{R}^{N\times n}.\end{split}$
Proof.
We first show the validity of the expression of $\boldsymbol{V}$ in (B.11). Defining $\boldsymbol{v}_{i}\mathrel{\mathrel{\mathop{:}}=}\boldsymbol{\Psi}^{\ast}\widehat{F}_{i}\in\mathbb{R}^{N}$, we consider the interpretation of $\boldsymbol{\Psi}^{\ast}$ as in the proof of Lemma B.5. Then, $\boldsymbol{v}_{i}=\left[\langle\widehat{F}_{i},\psi_{m}\rangle_{\mathcal{H}}\right]_{m=1}^{N}$. From the definition of $\widehat{F}_{i}$ in (4.3),
$\langle\widehat{F}_{i},\psi_{m}\rangle_{\mathcal{H}}=\langle\psi_{i}-\widehat{\mu},\psi_{m}\rangle_{\mathcal{H}}-\langle\psi_{i}-\widehat{\mu},\psi_{m}\rangle_{\mathcal{H}}\,\langle\psi_{i}-\widehat{\mu},\widehat{h}_{\lambda}\rangle_{\mathcal{H}}+\frac{1}{N}\sum_{i=1}^{N}\langle\widehat{\mu},\psi_{m}\rangle_{\mathcal{H}}\,\langle\psi_{i}-\widehat{\mu},\widehat{h}_{\lambda}\rangle_{\mathcal{H}}.$
The pairwise entries of $\widetilde{\boldsymbol{G}}$ and $\widetilde{\boldsymbol{h}}$ are respectively $\langle\psi_{i}-\widehat{\mu},\psi_{m}\rangle_{\mathcal{H}}$ and $\langle\psi_{i}-\widehat{\mu},\widehat{h}_{\lambda}\rangle_{\mathcal{H}}$, see Equations E.6 and E.8. Moreover, the column means are given the entries of the row vector $\frac{1}{N}\widetilde{\boldsymbol{h}}^{\top}\widetilde{\boldsymbol{G}}_{\mathcal{G}}$. Hence, the matrix form of $\boldsymbol{V}$ is precisely $\boldsymbol{V}=\left(\boldsymbol{I}-\operatorname{diag}(\widetilde{\boldsymbol{h}})\right)\widetilde{\boldsymbol{G}}+\frac{1}{N}\boldsymbol{1}\left(\widetilde{\boldsymbol{h}}^{\top}\widetilde{\boldsymbol{G}}\right)\in\mathbb{R}^{N\times N}$. Using Lemma B.5,
$\langle\widehat{F}_{i},\widehat{u}_{j}\rangle_{\mathcal{H}}=\frac{1}{\lambda}\langle\widehat{F}_{i},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}-\frac{1}{\lambda}\langle\boldsymbol{\Psi}^{\ast}\widehat{F}_{i},\boldsymbol{\gamma}_{j}\rangle_{\mathbb{R}^{N}}=\frac{1}{\lambda}\left(\beta_{i,j}-\boldsymbol{\gamma}_{j}^{\top}\boldsymbol{v}_{i}\right),$
where $\beta_{i,j}=\langle\widehat{F}_{i},\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})\rangle_{\mathcal{H}}$ is defined in the proof of Lemma B.6. Setting $\boldsymbol{S}\mathrel{\mathrel{\mathop{:}}=}\left[\langle\widehat{F}_{i},\widehat{u}_{j}\rangle_{\mathcal{H}}\right]_{i,j=1}^{N,n}\in\mathbb{R}^{N\times n}$, we obtain $\boldsymbol{S}=\frac{1}{\lambda}(\boldsymbol{B}-\boldsymbol{V}^{\top}\boldsymbol{\Lambda})$. Finally, from the definition of $\widehat{\boldsymbol{\Omega}}_{\lambda}$ in (4.5), we obtain $\widehat{\boldsymbol{\Omega}}_{\lambda}=\frac{1}{N}\boldsymbol{S}^{\top}\boldsymbol{S}\in\mathbb{R}^{n\times n}$. ∎
Appendix C Auxiliary lemmas for Section 4.1.3
Lemma C.1 (Continuity in the metric).
Let $\mathcal{M}\subset\mathbb{R}^{n}$ be a non-empty, closed, convex cone. Then for any sequence $\{\boldsymbol{M}_{k}\}_{k\in\mathbb{N}}\in\mathbb{S}^{n}_{++}$ converging to $\boldsymbol{M}\in\mathbb{S}^{n}_{++}$, it holds,
(C.1) $\underset{k\to\infty}{\lim}\,\,\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})=\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\quad\text{for any}\quad\boldsymbol{x}\in\mathbb{R}^{n}.$
Proof.
Fix any $\boldsymbol{x}\in\mathbb{R}^{n}$. From Weyl’s perturbation theorem, see Bhatia [1997, Corollary II.2.6], we obtain
$|\lambda_{j}(\boldsymbol{M}_{k})-\lambda_{j}(\boldsymbol{M})|\leq\|\boldsymbol{M}_{k}-\boldsymbol{M}\|_{\text{op}}\quad\text{for all}\quad j=1,\ldots,n.$
Since $\|\boldsymbol{M}_{k}-\boldsymbol{M}\|_{\text{op}}\to 0$ as $k\to\infty$, hence $\lambda_{j}(\boldsymbol{M}_{k})\to\lambda_{j}(\boldsymbol{M})$ as $k\to\infty$ for $1\leq j\leq n$. Thus, the sequence of eigenvalues of $\boldsymbol{M}_{k}$ is bounded, i.e., there exist real numbers $c,C>0$ such that:
(C.2) $0<c\leq\lambda_{\min}(\boldsymbol{M}_{k})\leq\lambda_{\max}(\boldsymbol{M}_{k})\leq C<\infty\quad\text{for all}\quad k\in\mathbb{N}.$
The above inequality implies that
(C.3) $c\|\boldsymbol{x}\|_{2}^{2}\leq\boldsymbol{x}^{\top}\boldsymbol{M}_{k}\,\boldsymbol{x}=\|\boldsymbol{x}\|_{\boldsymbol{M}_{k}}^{2}\leq C\|\boldsymbol{x}\|_{2}^{2}\quad\text{for all}\quad k\in\mathbb{N}.$
Now, for any $k\in\mathbb{N}$, consider the Hilbert space $\mathbb{R}^{n}$ equipped with the inner product $\langle\cdot,\cdot\rangle_{\boldsymbol{M}_{k}}$, and set $\boldsymbol{u}_{k}\mathrel{\mathrel{\mathop{:}}=}\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})$. From the best approximation property of a projection, see Bauschke and Combettes [2017, Chapter 3.2], it follows that
$\|\boldsymbol{u}_{k}-\boldsymbol{x}\|_{\boldsymbol{M}_{k}}\leq\|\boldsymbol{u}-\boldsymbol{x}\|_{\boldsymbol{M}_{k}}\quad\text{for any}\quad\boldsymbol{u}\in\mathcal{M}.$
Choosing $\boldsymbol{u}=\boldsymbol{0}$ in the above inequality and using (C.3) gives
$\|\boldsymbol{u}_{k}\|_{\boldsymbol{M}_{k}}\leq\|\boldsymbol{x}\|_{\boldsymbol{M}_{k}}+\|\boldsymbol{u}_{k}-\boldsymbol{x}\|_{\boldsymbol{M}_{k}}\leq 2\|\boldsymbol{x}\|_{\boldsymbol{M}_{k}}\leq 2\sqrt{C}\|\boldsymbol{x}\|_{2}\quad\text{for all }k\in\mathbb{N}.$
Using (C.3) again and the above inequality, we have
$\|\boldsymbol{u}_{k}\|_{2}\leq\frac{1}{\sqrt{c}}\|\boldsymbol{u}_{k}\|_{\boldsymbol{M}_{k}}\leq 2\sqrt{\frac{C}{c}}\,\|\boldsymbol{x}\|_{2}\quad\text{for all }k\in\mathbb{N}.$
Thus, $\boldsymbol{u}_{k}$ is bounded in the Euclidean norm and hence, by the Bolzano-Weierstrass theorem, there exists a convergent subsequence $\boldsymbol{u}_{k_{\ell}}\to{\boldsymbol{u}}^{\ast}$, where we consider the convergence in the topology induced by the usual Euclidean norm. From Theorem 4.5, the projection $\boldsymbol{u}_{k_{\ell}}$ uniquely satisfies
$\Bigl\langle\boldsymbol{x}-\boldsymbol{u}_{k_{\ell}},\boldsymbol{v}-\boldsymbol{u}_{k_{\ell}}\Bigr\rangle_{\boldsymbol{M}_{k_{\ell}}}=\Bigl\langle\boldsymbol{x}-\boldsymbol{u}_{k_{\ell}},\boldsymbol{M}_{k_{\ell}}\left(\boldsymbol{v}-\boldsymbol{u}_{k_{\ell}}\right)\Bigr\rangle_{\mathbb{R}^{n}}\leq 0\quad\text{for any }\boldsymbol{v}\in\mathcal{M}.$
We can split the expression into two terms as
$\underbrace{\Bigl\langle\boldsymbol{x}-\boldsymbol{u}_{k_{\ell}},\boldsymbol{M}\left(\boldsymbol{v}-\boldsymbol{u}_{k_{\ell}}\right)\Bigr\rangle_{\mathbb{R}^{n}}}_{(I)}+\underbrace{\Bigl\langle\boldsymbol{x}-\boldsymbol{u}_{k_{\ell}},\left(\boldsymbol{M}_{k_{\ell}}-\boldsymbol{M}\right)\left(\boldsymbol{v}-\boldsymbol{u}_{k_{\ell}}\right)\Bigr\rangle_{\mathbb{R}^{n}}}_{(II)}.$
By the continuity of the bilinear form induced by the Euclidean inner product,
$(I)\to\Bigl\langle\boldsymbol{x}-\boldsymbol{u}^{\ast},\boldsymbol{M}(\boldsymbol{v}-\boldsymbol{u}^{\ast})\Bigr\rangle_{\mathbb{R}^{n}}\quad\text{as}\quad\ell\to\infty.$
Now, for any fixed $\boldsymbol{x}\in\mathbb{R}^{n},\,\boldsymbol{v}\in\mathcal{M}$, the terms $(\boldsymbol{x}-\boldsymbol{u}_{k_{\ell}}),\,(\boldsymbol{v}-\boldsymbol{u}_{k_{\ell}})$ are bounded since $\boldsymbol{u}_{k_{\ell}}$ is a convergent sequence. Hence,
$\Bigl|\Bigl\langle\boldsymbol{x}-\boldsymbol{u}_{k_{\ell}},\left(\boldsymbol{M}_{k_{\ell}}-\boldsymbol{M}\right)\left(\boldsymbol{v}-\boldsymbol{u}_{k_{\ell}}\right)\Bigr\rangle_{\mathbb{R}^{n}}\Bigr|\leq\|\boldsymbol{M}_{k_{\ell}}-\boldsymbol{M}\|_{\text{op}}\,\|\boldsymbol{x}-\boldsymbol{u}_{k_{\ell}}\|_{2}\,\|\boldsymbol{v}-\boldsymbol{u}_{k_{\ell}}\|_{2}.$
Taking the limit as $\ell\to\infty$, we have $(II)\to 0$. Hence, for any $\boldsymbol{v}\in\mathcal{K}$, it holds,
$\underset{\ell\to\infty}{\lim}\,\,\Bigl\langle\boldsymbol{x}-\boldsymbol{u}_{k_{\ell}},\boldsymbol{v}-\boldsymbol{u}_{k_{\ell}}\Bigr\rangle_{\boldsymbol{M}_{k_{\ell}}}=\Bigl\langle\boldsymbol{x}-\boldsymbol{u}^{\ast},\boldsymbol{M}(\boldsymbol{v}-\boldsymbol{u}^{\ast})\Bigr\rangle_{\mathbb{R}^{n}}=\langle\boldsymbol{x}-\boldsymbol{u}^{\ast},\boldsymbol{v}-\boldsymbol{u}^{\ast}\rangle_{\boldsymbol{M}}\leq 0.$
But this is the inequality characterizing the unique projection $\boldsymbol{u}^{\ast}=\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})$, see Theorem 4.5. Hence, the set of subsequential limits of $\boldsymbol{u}_{k}$ is unique and any convergent subsequence of $\{\boldsymbol{u}_{k}\}_{k\in\mathbb{N}}$ has the same limit $\boldsymbol{u}^{\ast}$. This property and the fact that $\{\boldsymbol{u}_{k}\}$ is bounded in the Euclidean norm imply that the sequence $\boldsymbol{u}_{k}$ converges to the limit $\boldsymbol{u}^{\ast}$. Since $\boldsymbol{x}\in\mathbb{R}^{n}$ is arbitrary, therefore,
$\underset{k\to\infty}{\lim}\,\,\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})=\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\quad\text{for any }\boldsymbol{x}\in\mathbb{R}^{n},$
where the convergence is in the usual topology on $\mathbb{R}^{n}$ generated by the Euclidean norm. ∎
Lemma C.2 (Joint continuity of projection).
Let $\mathcal{M}\subset\mathbb{R}^{n}$ be a non-empty, closed, convex cone. The map $f\colon\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}\to\mathbb{R}^{n}$ defined as
(C.4) $f(\boldsymbol{x},\boldsymbol{M})\mathrel{\mathrel{\mathop{:}}=}\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\quad\text{for any }(\boldsymbol{x},\boldsymbol{M})\in\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}$
is jointly continuous.
Proof.
Consider any sequence $(\boldsymbol{x}_{k},\boldsymbol{M}_{k})\in\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}$ that converges to some fixed $(\boldsymbol{x},\boldsymbol{M})\in\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}$ in the usual product topology generated by the respective norms. Therefore,
(C.5) $\|\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x}_{k})-\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\|_{2}\leq\|\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x}_{k})-\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})\|_{2}+\|\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})-\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\|_{2}.$
Since $\mathcal{M}$ is a non-empty, closed, convex cone, the projection $\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\cdot)$ is Lipschitz, see Bauschke and Combettes [2017, Definition 4.1 and Proposition 4.16]. Hence,
$\|\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x}_{k})-\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})\|_{\boldsymbol{M}_{k}}\leq\|\boldsymbol{x}_{k}-\boldsymbol{x}\|_{\boldsymbol{M}_{k}}\quad\text{for any }k\in\mathbb{N}.$
Using (C.3) in the above, we obtain that for all $k\in\mathbb{N}$,
$\|\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x}_{k})-\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})\|_{2}\leq\frac{1}{\sqrt{c}}\|\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x}_{k})-\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})\|_{\boldsymbol{M}_{k}}\leq\frac{1}{\sqrt{c}}\|\boldsymbol{x}_{k}-\boldsymbol{x}\|_{\boldsymbol{M}_{k}}\leq\sqrt{\frac{C}{c}}\|\boldsymbol{x}_{k}-\boldsymbol{x}\|_{2}.$
Taking the limit $k\to\infty$, it follows that $\|\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x}_{k})-\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})\|_{2}\to 0$ since $\|\boldsymbol{x}_{k}-\boldsymbol{x}\|_{2}\to 0$. Finally, from Lemma C.1, $\|\Pi^{\boldsymbol{M}_{k}}_{\mathcal{M}}(\boldsymbol{x})-\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\|_{2}\to 0$ as $k\to\infty$. Hence, the function $f$ mapping $(\boldsymbol{x},\boldsymbol{M})\mapsto\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})$ is jointly continuous on $\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}$. ∎
Lemma C.3 (Continuity of squared projection error).
Let $\mathcal{M}\subset\mathbb{R}^{n}$ be a non-empty, closed, convex cone. The map $g\colon\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}\to\mathbb{R}$ defined as
(C.6) $g(\boldsymbol{x},\boldsymbol{M})\mathrel{\mathrel{\mathop{:}}=}\|\boldsymbol{x}-\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\|_{\boldsymbol{M}}^{2}\quad\text{for any }(\boldsymbol{x},\boldsymbol{M})\in\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}$
is jointly continuous.
Proof.
Define the following maps
$\begin{split}f_{1}(\boldsymbol{x},\boldsymbol{M})&\mathrel{\mathrel{\mathop{:}}=}(\boldsymbol{x},\boldsymbol{M},\boldsymbol{\Pi}^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})),\qquad f_{2}(\boldsymbol{x},\boldsymbol{M},\boldsymbol{y})\mathrel{\mathrel{\mathop{:}}=}\langle\boldsymbol{x}-\boldsymbol{y},\boldsymbol{M}(\boldsymbol{x}-\boldsymbol{y})\rangle_{\mathbb{R}^{n}}.\end{split}$
From Lemma C.2, we can conclude that $f_{1}$ is continuous, while $f_{2}$ is continuous from the continuity of the bilinear form induced by the Euclidean inner product on $\mathbb{R}^{n}$. Hence,
$g(\boldsymbol{x},\boldsymbol{M})=\|\boldsymbol{x}-\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\|_{\boldsymbol{M}}^{2}=\Bigl\langle\boldsymbol{x}-\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x}),\boldsymbol{M}\big(\boldsymbol{x}-\Pi^{\boldsymbol{M}}_{\mathcal{M}}(\boldsymbol{x})\big)\Bigr\rangle_{\mathbb{R}^{n}}=f_{2}\circ f_{1}(\boldsymbol{x},\boldsymbol{M})$
is jointly continuous in its arguments as a composition of continuous maps. ∎
Proof of Theorem 4.6.
By the continuity of the inversion operation on $\mathbb{S}^{n}_{++}$, we have from the consistency of $\widehat{\boldsymbol{\Omega}}_{\lambda}$ in Theorem 4.2 that
(C.7) $\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}\overset{a.s.}{\longrightarrow}\boldsymbol{\Omega}_{\lambda}^{-1}\implies\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}\overset{\mathbb{P}}{\longrightarrow}\boldsymbol{\Omega}_{\lambda}^{-1}\quad\text{as}\quad N\to\infty.$
Now, define $\boldsymbol{Z}_{N}\mathrel{\mathrel{\mathop{:}}=}\sqrt{N}\widehat{\boldsymbol{\theta}}$. Under the least favorable null $H_{0}:\boldsymbol{\theta}=\boldsymbol{0}$, we have from Proposition 4.1,
$\boldsymbol{Z}_{N}\overset{d}{\longrightarrow}\boldsymbol{Z}\sim\mathcal{N}_{n}(\boldsymbol{0},\boldsymbol{\Omega}_{\lambda}).$
Hence, we have from van der Vaart [1998, Theorem 2.7], for asymptotically large $N$,
$\bigl(\boldsymbol{Z}_{N},\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}\bigr)\overset{d}{\longrightarrow}\bigl(\boldsymbol{Z},\boldsymbol{\Omega}_{\lambda}^{-1}\bigr)\quad\text{on}\quad\mathbb{R}^{n}\times\mathbb{S}^{n}_{++}.$
Since $\mathbb{R}^{n}_{+}$ is a closed convex cone, $\sqrt{N}\mathbb{R}^{n}_{+}=\mathbb{R}^{n}_{+}$ for any $N\geq 1$. Hence,
$W_{N}$ $=\underset{\boldsymbol{c}\in\mathbb{R}^{n}_{+}}{\min}\,\,N(\widehat{\boldsymbol{\theta}}-\boldsymbol{c})^{\top}\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}(\widehat{\boldsymbol{\theta}}-\boldsymbol{c})$
$=\underset{\boldsymbol{c}\in\mathbb{R}^{n}_{+}}{\min}\,\,(\sqrt{N}\widehat{\boldsymbol{\theta}}-\sqrt{N}\boldsymbol{c})^{\top}\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}(\sqrt{N}\widehat{\boldsymbol{\theta}}-\sqrt{N}\boldsymbol{c})$
$=\underset{\boldsymbol{u}\in\mathbb{R}^{n}_{+}}{\min}\,\,(\boldsymbol{Z}_{N}-\boldsymbol{u})^{\top}\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}(\boldsymbol{Z}_{N}-\boldsymbol{u})$
$=\|\boldsymbol{Z}_{N}-\Pi^{\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}}_{\mathbb{R}^{n}_{+}}(\boldsymbol{Z}_{N})\|_{\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}}^{2}.$
From Lemma C.3, $W_{N}$ is continuous as a function of $\bigl(\boldsymbol{Z}_{N},\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}\bigr)$. Hence, by the CMT,
$W_{N}\overset{d}{\longrightarrow}W\mathrel{\mathrel{\mathop{:}}=}\|\boldsymbol{Z}-\Pi^{\boldsymbol{\Omega}_{\lambda}^{-1}}_{\mathbb{R}^{n}_{+}}(\boldsymbol{Z})\|^{2}_{{\boldsymbol{\Omega}}_{\lambda}^{-1}}.$
From (4.12), it follows, $W\sim\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\big(\mathbb{R}^{n}_{+}\bigr)^{\circ})$. Moreover, from Moreau’s decomposition and the Pythagorean identity,
$\boldsymbol{Z}^{\top}{\boldsymbol{\Omega}}_{\lambda}^{-1}\boldsymbol{Z}=\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+})+\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\big(\mathbb{R}^{n}_{+}\bigr)^{\circ})$
Since $\boldsymbol{Z}\sim\mathcal{N}_{n}(\boldsymbol{0},\boldsymbol{\Omega}_{\lambda})$, hence $\boldsymbol{Z}^{\top}\boldsymbol{\Omega}_{\lambda}^{-1}\boldsymbol{Z}\sim\chi^{2}_{n}$. Thus, $\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\big(\mathbb{R}^{n}_{+}\bigr)^{\circ})=\chi^{2}_{n}-\bar{\chi}^{2}(\boldsymbol{\Omega}_{\lambda},\mathbb{R}^{n}_{+})$, where the equality holds almost surely. ∎
Appendix D Implementation
In this section, derivatives are indexed by the complete set $\mathcal{A}_{s}=\{\boldsymbol{\alpha}:\lvert\boldsymbol{\alpha}\rvert\leq s\}$. For cases where a subset $\mathcal{A}\subset\mathcal{A}_{s}$ is employed, we set $w_{\boldsymbol{\alpha}}\equiv 0$ for any $\boldsymbol{\alpha}\notin\mathcal{A}$; thus, all subsequent statements and formulations remain unchanged.
D.1. Matrix formulation
For any $\boldsymbol{\alpha}\in\mathcal{A}_{s}$, define the row vectors of basis functions
(D.1) $\boldsymbol{\Phi}^{(\boldsymbol{\alpha})}\mathrel{\mathrel{\mathop{:}}=}[\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x}_{1}),\ldots,\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x}_{N})].$
Denote the canonical basis vectors of $\mathbb{R}^{d}$ by $\{\boldsymbol{e}_{j}\}_{j=1}^{d}$ and consider the following ordering of $\mathcal{A}_{s}$:
(D.2) $\Bigl[1,\alpha_{1},\ldots,\alpha_{d},\alpha_{1}^{2},\alpha_{1}\alpha_{2},\ldots,\alpha_{d}^{2},\ldots,\alpha_{1}^{s},\ldots,\alpha_{d}^{s}\Bigr].$
We stack basis functions of the optimal subspace $\mathcal{H}_{X}$ as the row vector:
(D.3) $\boldsymbol{\Phi}\mathrel{\mathrel{\mathop{:}}=}\Big[\boldsymbol{\Phi}^{(\boldsymbol{\alpha})}\Big]_{\boldsymbol{\alpha}\in\mathcal{A}_{s}},$
where we consider the ordering as in (D.2). So, $\boldsymbol{\Phi}$ has $M\mathrel{\mathrel{\mathop{:}}=}Nm_{s}$ columns. We now define the corresponding kernel matrix by taking the pairwise inner product:
(D.4) $\boldsymbol{K}\mathrel{\mathrel{\mathop{:}}=}\langle\boldsymbol{\Phi}^{\top},\boldsymbol{\Phi}\rangle_{\mathcal{H}}\in\mathbb{R}^{M\times M}.$
Since we have the functional form of $\widehat{h}_{\lambda}$ as in (2.13), we can now write:
(D.5) $\widehat{h}_{\lambda}=\boldsymbol{\Phi}\widehat{\boldsymbol{c}},\qquad\widehat{\boldsymbol{c}}\mathrel{\mathrel{\mathop{:}}=}\bigl[\widehat{c}_{i,\boldsymbol{\alpha}}\bigr]\in\mathbb{R}^{M}.$
Note that we follow the same ordering that is compatible with the ordering of the basis functions in $\boldsymbol{\Phi}$. We need to formulate the system of equations that solves for the optimal coefficients $\widehat{\boldsymbol{c}}$. Hence, we now proceed to write each term in Problem 2.12 in terms of the matrix formulation which will lead to the desired system of equations.
Towards that end, we seek to write $\psi_{i}$ from (2.7) in terms of $\boldsymbol{\Phi}$. Consider the following construction: for any $\boldsymbol{\alpha}\in\mathcal{A}_{s}$, define
(D.6) $\boldsymbol{A}^{(\boldsymbol{\alpha})}\mathrel{\mathrel{\mathop{:}}=}\operatorname{diag}\left(w_{\boldsymbol{\alpha}}(\boldsymbol{z}_{1}),\ldots,w_{\boldsymbol{\alpha}}(\boldsymbol{z}_{N})\right)\in\mathbb{R}^{N\times N}.$
We now define the block matrix of coefficients:
(D.7) $\boldsymbol{A}\mathrel{\mathrel{\mathop{:}}=}\begin{bmatrix}\boldsymbol{A}^{(\boldsymbol{\alpha}_{1})}\\ \vdots\\ \boldsymbol{A}^{(\boldsymbol{\alpha}_{m_{s}})}\end{bmatrix}\in\mathbb{R}^{M\times N},\qquad M=Nm_{s},$
where we use the same ordering as in (D.2). Define the column vectors $\boldsymbol{a}_{i}\mathrel{\mathrel{\mathop{:}}=}[\boldsymbol{A}_{:,i}]\in\mathbb{R}^{M}$ for $1\leq i\leq n$, that satisfies
(D.8) $\psi_{i}=\sum_{\boldsymbol{\alpha}\mathcal{A}_{s}}w_{\boldsymbol{\alpha}}(\boldsymbol{z}_{i})\phi^{(\boldsymbol{\alpha})}(\boldsymbol{x}_{i})=\boldsymbol{\Phi}\boldsymbol{a}_{i}.$
Consider the mean vector
$\bar{\boldsymbol{a}}\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathbb{E}}[\boldsymbol{a}_{i}]=\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{a}_{i}\in\mathbb{R}^{M},$
and the centered vectors
$\widetilde{\boldsymbol{a}}_{i}\mathrel{\mathrel{\mathop{:}}=}\boldsymbol{a}_{i}-\bar{\boldsymbol{a}}\quad\text{for}\quad 1\leq i\leq N.$
Hence, we can write $\widehat{\mu}=\widehat{\mathbb{E}}[\psi_{i}]=\widehat{\mathbb{E}}[\boldsymbol{\Phi}\boldsymbol{a}_{i}]=\boldsymbol{\Phi}\bar{\boldsymbol{a}}$ such that:
$\langle h,\widehat{\mu}\rangle_{\mathcal{H}}=\langle\boldsymbol{c}^{\top}\boldsymbol{\Phi}^{\top},\boldsymbol{\Phi}\bar{\boldsymbol{a}}\rangle_{\mathcal{H}}=\boldsymbol{c}^{\top}\boldsymbol{K}\bar{\boldsymbol{a}}.$
Now, it holds:
$\langle h,\psi_{i}-\widehat{\mu}\rangle_{\mathcal{H}}=\langle\boldsymbol{c}^{\top}\boldsymbol{\Phi}^{\top},\boldsymbol{\Phi}\widetilde{\boldsymbol{a}}_{i}\rangle_{\mathcal{H}}=\boldsymbol{c}^{\top}\boldsymbol{K}\widetilde{\boldsymbol{a}}_{i}.$
Therefore, the variance term reads
$\widehat{\mathbb{E}}[\langle h,\psi_{i}-\widehat{\mu}\rangle_{\mathcal{H}}^{2}]=\widehat{\mathbb{E}}\left[(\boldsymbol{c}^{\top}\boldsymbol{K}\widetilde{\boldsymbol{a}}_{i})^{2}\right]=\boldsymbol{c}^{\top}\boldsymbol{K}\boldsymbol{\Sigma}\boldsymbol{K}\boldsymbol{c},$
where
$\boldsymbol{\Sigma}\mathrel{\mathrel{\mathop{:}}=}\widehat{\mathbb{E}}\left[\widetilde{\boldsymbol{a}}_{i}\widetilde{\boldsymbol{a}}_{i}^{\top}\right]=\frac{1}{N}\sum_{i=1}^{N}\boldsymbol{a}_{i}\boldsymbol{a}_{i}^{\top}.$
Finally, the regularization term can be written as
$\langle h,h\rangle_{\mathcal{H}}=\langle\boldsymbol{c}^{\top}\boldsymbol{\Phi}^{\top},\boldsymbol{\Phi}\boldsymbol{c}\rangle_{\mathcal{H}}=\boldsymbol{c}^{\top}\boldsymbol{K}\boldsymbol{c}.$
Having computed the above terms, the matrix formulation of Problem 2.12 is given exactly by Problem D.9 below:
(D.9) $\begin{split}\widehat{\boldsymbol{c}}=\,\,\underset{\boldsymbol{c}\in\mathbb{R}^{M}}{\operatorname{argmin}}\,\,-\boldsymbol{c}^{\top}\boldsymbol{K}\bar{\boldsymbol{a}}+\frac{1}{2}\boldsymbol{c}^{\top}\boldsymbol{K}\boldsymbol{\Sigma}\boldsymbol{K}\boldsymbol{c}+\frac{\lambda}{2}\boldsymbol{c}^{\top}\boldsymbol{K}\boldsymbol{c}.\end{split}$
D.2. Efficient computation
Problem D.9 is convex and has a unique minimum, which, from the first-order conditions, can be obtained as
(D.10) $(\boldsymbol{K}\boldsymbol{\Sigma}\boldsymbol{K}+\lambda\boldsymbol{K})\,\widehat{\boldsymbol{c}}=\boldsymbol{K}\,\bar{\boldsymbol{a}}.$
Solving (D.10) for $\widehat{\boldsymbol{c}}$ can be computationally demanding and memory-intensive, especially when we have a large number of observations/large dimension/large number of derivative evaluations, since $M$ depends on $N,d,s$. In particular, the computational cost of solving (D.10) in a naive way is cubic $\mathcal{O}(M^{3})$, while the formation and storage of the full kernel matrix $\boldsymbol{K}$ is $\mathcal{O}(M^{2})$ in memory. As a result, we need an efficient way to solve (D.10) such that we can lessen our computational and storage requirements. This is facilitated by pivoted Cholesky decomposition of Harbrecht et al. [2012]. We show here how to leverage the pivoted Cholesky of $\boldsymbol{K}$ to reduce the computational burden. In addition, we remark that this algorithm does not necessitate forming the full kernel matrix and thus also helps in reducing the storage cost.
We first consider the pivoted Cholesky decomposition of $\boldsymbol{K}$ as:
$\boldsymbol{K}\approx\boldsymbol{L}\boldsymbol{L}^{\top},\qquad\boldsymbol{L}\in\mathbb{R}^{M\times m},\quad m\ll M.$
From this algorithm, we have the following relations between the biorthogonal matrix and the Cholesky factor, see Filipović et al. [2025, Theorem 4.1]
$\boldsymbol{K}\boldsymbol{B}=\boldsymbol{L},\qquad\boldsymbol{B}^{\top}\boldsymbol{L}=\boldsymbol{L}^{\top}\boldsymbol{B}=\boldsymbol{I}_{m},\qquad\boldsymbol{B}\in\mathbb{R}^{M\times m}.$
Premultiplying both sides of (D.10) with $\boldsymbol{B}^{\top}$,
$(\boldsymbol{L}^{\top}\boldsymbol{\Sigma}\boldsymbol{L}\boldsymbol{L}^{\top}+\lambda\boldsymbol{L}^{\top})\,\widehat{\boldsymbol{c}}=\boldsymbol{L}^{\top}\bar{\boldsymbol{a}}.$
Next, we define the vectors
$\boldsymbol{L}^{\top}\,\widehat{\boldsymbol{c}}=\widetilde{\boldsymbol{c}},\qquad\boldsymbol{L}^{\top}\bar{\boldsymbol{a}}=\widetilde{\boldsymbol{b}}$.
Hence, we can now write:
(D.11) $(\boldsymbol{L}^{\top}\boldsymbol{\Sigma}\boldsymbol{L}+\lambda\boldsymbol{I})\,\widetilde{\boldsymbol{c}}=\widetilde{\boldsymbol{b}}$.
We solve the above equation for $\widetilde{\boldsymbol{c}}$ and then use $\widehat{\boldsymbol{c}}=\boldsymbol{B}\,\widetilde{\boldsymbol{c}}$ to get back $\widehat{\boldsymbol{c}}$.
Remark D.1.
Solving (D.11) costs only $\mathcal{O}(m^{3})$, which is considerably cheaper as opposed to $\mathcal{O}(M^{3})$, since $m\ll M$. The columns of $\boldsymbol{B}$ span the same rank $m$-subspace as $\operatorname{Im}(\boldsymbol{L})$, and $\boldsymbol{B}$ acts as the left-inverse of $\boldsymbol{L}$. Hence, the matrix $\boldsymbol{B}\boldsymbol{L}^{\top}\in\mathbb{R}^{M\times M}$ acts the orthogonal projector onto $\operatorname{Im}(L)$. Since $\boldsymbol{c}\in\operatorname{Im}(\boldsymbol{L})$, pre-multiplying by $\boldsymbol{B}^{\top}$ shrinks (D.10) to $m$ dimensions; afterwards using $\boldsymbol{B}\widetilde{\boldsymbol{c}}$ to retrieve $\widehat{\boldsymbol{c}}$ gives the best-possible approximation within the low-rank subspace generated by the columns of the pivoted Cholesky factor $\boldsymbol{L}$.
Appendix E Construction of test statistic
In this section, we exhibit how to construct the test statistic to test the shape constraints of $h_{\lambda}$ jointly on the finite grid $\mathcal{G}=\{\boldsymbol{\xi}_{j}:1\leq j\leq n\}$, leveraging the sign of the derivative evaluation. We first define the row vector of functions
(E.1) $\boldsymbol{\Phi}_{\mathcal{G}}\mathrel{\mathrel{\mathop{:}}=}[\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{1}),\ldots,\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{n})],$
and the corresponding kernel matrix for the test grid
(E.2) $\boldsymbol{K}_{\mathcal{G}}\mathrel{\mathrel{\mathop{:}}=}\langle\boldsymbol{\Phi}^{\top},\boldsymbol{\Phi}_{\mathcal{G}}\rangle_{\mathcal{H}}\in\mathbb{R}^{M\times n},$
where $\boldsymbol{\Phi}$ is defined in (D.3). We write the row vector consisting of $\psi_{i}$ for $1\leq i\leq N$ as:
(E.3) $\boldsymbol{\Psi}\mathrel{\mathrel{\mathop{:}}=}[\psi_{1},\ldots,\psi_{N}]=\boldsymbol{\Phi}[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{N}]=\boldsymbol{\Phi}\boldsymbol{A},$
where $\boldsymbol{A}=[\boldsymbol{a}_{1},\ldots,\boldsymbol{a}_{N}]\in\mathbb{R}^{M\times N}$ is the matrix from (D.7), whose columns are given by $\boldsymbol{a}_{i}$, see Appendix D for more details. The corresponding Gram matrix (in the $\boldsymbol{\Psi}$ basis) is:
(E.4) $\boldsymbol{G}\mathrel{\mathrel{\mathop{:}}=}\langle\boldsymbol{\Psi}^{\top},\boldsymbol{\Psi}\rangle_{\mathcal{H}}=\boldsymbol{A}^{\top}\langle\boldsymbol{\Phi}^{\top},\boldsymbol{\Phi}\rangle_{\mathcal{H}}\boldsymbol{A}=\boldsymbol{A}^{\top}\boldsymbol{K}\boldsymbol{A}\in\mathbb{R}^{N\times N}.$
Define the centering matrix $\boldsymbol{H}\mathrel{\mathrel{\mathop{:}}=}\boldsymbol{I}_{N}-\frac{1}{N}\boldsymbol{1}\boldsymbol{1}^{\top}\in\mathbb{R}^{N\times N}$, where we define the column vector of ones $\boldsymbol{1}\mathrel{\mathrel{\mathop{:}}=}[1,\ldots,1]^{\top}\in\mathbb{R}^{N}$. Note that the matrix $\boldsymbol{H}$ is symmetric and idempotent, i.e., $\boldsymbol{H}^{\top}=\boldsymbol{H}$. We can define the following row vector of centered functions
(E.5) $\widetilde{\boldsymbol{\Psi}}\mathrel{\mathrel{\mathop{:}}=}[\psi_{1}-\widehat{\mu},\ldots,\psi_{N}-\widehat{\mu}]=\boldsymbol{\Psi}\boldsymbol{H},$
and the matrix
(E.6) $\widetilde{\boldsymbol{G}}\mathrel{\mathrel{\mathop{:}}=}\langle\widetilde{\boldsymbol{\Psi}}^{\top},\boldsymbol{\Psi}\rangle_{\mathcal{H}}=\boldsymbol{H}\boldsymbol{G}.$
Now, we define the Gram matrix with respect to the centered basis functions as
(E.7) $\widetilde{\boldsymbol{G}}_{\mathcal{G}}\mathrel{\mathrel{\mathop{:}}=}\langle\widetilde{\boldsymbol{\Psi}}^{\top},\boldsymbol{\Phi}_{\mathcal{G}}\rangle_{\mathcal{H}}=\boldsymbol{H}\langle\boldsymbol{\Psi}^{\top},\boldsymbol{\Phi}_{\mathcal{G}}\rangle_{\mathcal{H}}=\boldsymbol{H}\boldsymbol{A}^{\top}\langle\boldsymbol{\Phi}^{\top},\boldsymbol{\Phi}_{\mathcal{G}}\rangle_{\mathcal{H}}=\boldsymbol{H}\boldsymbol{A}^{\top}\boldsymbol{K}_{\mathcal{G}}\in\mathbb{R}^{N\times n},$
and the sample-estimator $\widehat{h}_{\lambda}$ (in this basis) as
(E.8) $\widetilde{\boldsymbol{h}}\mathrel{\mathrel{\mathop{:}}=}\left[\langle\psi_{i}-\widehat{\mu},\widehat{h}_{\lambda}\rangle_{\mathcal{H}}\right]_{i=1}^{N}=\langle\widetilde{\boldsymbol{\Psi}}^{\top},\boldsymbol{\Phi}\rangle_{\mathcal{H}}\,\widehat{\boldsymbol{c}}=\boldsymbol{H}\boldsymbol{A}^{\top}\langle\boldsymbol{\Phi}^{\top},\boldsymbol{\Phi}\rangle_{\mathcal{H}}\,\widehat{\boldsymbol{c}}=\boldsymbol{H}\boldsymbol{A}^{\top}\boldsymbol{K}\widehat{\boldsymbol{c}}\in\mathbb{R}^{N}.$
From Lemma B.4, the action of the sample covariance operator $\widehat{\mathnormal{\Sigma}}$, cp. (2.10) may be realized as $\frac{1}{N}\widetilde{\boldsymbol{\Psi}}\widetilde{\boldsymbol{\Psi}}^{\ast}$ and thus, we can construct the sample analogue of $u_{j}$, that is, $\widehat{u}_{j}$ from (4.2) as
(E.9) $\widehat{u}_{j}=\widehat{\mathnormal{\Sigma}}_{\lambda}^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})=\left(\frac{1}{N}\widetilde{\boldsymbol{\Psi}}\widetilde{\boldsymbol{\Psi}}^{\ast}+\lambda I\right)^{-1}\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j}),$
which can be computed as $\widehat{u}_{j}=\frac{1}{\lambda}(\phi^{(\boldsymbol{\alpha})}(\boldsymbol{\xi}_{j})-\boldsymbol{\Psi}\boldsymbol{\gamma}_{j})$, see Lemma B.5, where
(E.10) $\boldsymbol{\gamma}_{j}\mathrel{\mathrel{\mathop{:}}=}\frac{1}{N}\boldsymbol{H}\left(\lambda\boldsymbol{I}_{N}+\frac{1}{N}\boldsymbol{H}\boldsymbol{G}\boldsymbol{H}\right)^{-1}[\widetilde{\boldsymbol{G}}_{\mathcal{G}}]_{j}\in\mathbb{R}^{N}.$
Now, Lemma B.7 directly gives us a computational solution for constructing the finite-sample covariance estimator matrix $\widehat{\boldsymbol{\Omega}}_{\lambda}$ in closed-form. Having computed $\widehat{\boldsymbol{\Omega}}_{\lambda}$, we proceed to compute the test statistic $W_{N}$ from Theorem 4.6 as follows. Consider the vector stacked evaluations of the derivative functional at the grid points $\widehat{\boldsymbol{\theta}}$. Set $\boldsymbol{b}\mathrel{\mathrel{\mathop{:}}=}\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1/2}\,\widehat{\boldsymbol{\theta}}$, where $\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1/2}$ is a matrix root of $\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1}$. Then, we can write:
$W_{N}\mathrel{\mathrel{\mathop{:}}=}N\,\min_{\boldsymbol{c}\in\mathbb{R}^{n}_{+}}\,\,(\widehat{\boldsymbol{\theta}}-\boldsymbol{c})^{\top}\widehat{\Omega}_{\lambda}^{-1}(\widehat{\boldsymbol{\theta}}-\boldsymbol{c})=N\,\min_{\boldsymbol{c}\in\mathbb{R}^{n}_{+}}\,\,\|\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1/2}\boldsymbol{c}-\boldsymbol{b}\|_{2}^{2}.$
The optimization problem has a unique minimizer
$\boldsymbol{c}^{\star}\mathrel{\mathrel{\mathop{:}}=}\min_{\boldsymbol{c}\in\mathbb{R}^{n}_{+}}\,\,\|\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1/2}\boldsymbol{c}-\boldsymbol{b}\|_{2}^{2},$
that can be solved as a non-negative least-squares program. Define the residuals $\boldsymbol{r}\mathrel{\mathrel{\mathop{:}}=}\widehat{\boldsymbol{\Omega}}_{\lambda}^{-1/2}\boldsymbol{c}^{\star}-\boldsymbol{b}$. Then, we can compute the test statistic as $W_{N}=N\|\boldsymbol{r}\|_{2}^{2}$.